### v1

# DreamTalk: When Expressive Talking Head Generation Meets Diffusion Probabilistic Models

Abstract
Diffusion models have shown remarkable success in a variety of downstream generative tasks, yet remain under-explored in the important and challenging expressive talking head generation. In this work, we propose a DreamTalk framework to fulfill this gap, which employs meticulous design to unlock the potential of diffusion models in generating expressive talking heads. Specifically, DreamTalk consists of three crucial components: a denoising network, a style-aware lip expert, and a style predictor. The diffusion-based denoising network is able to consistently synthesize high-quality audio-driven face motions across diverse expressions. To enhance the expressiveness and accuracy of lip motions, we introduce a style-aware lip expert that can guide lip-sync while being mindful of the speaking styles. To eliminate the need for expression reference video or text, an extra diffusion-based style predictor is utilized to predict the target expression directly from the audio. By this means, DreamTalk can harness powerful diffusion models to generate expressive faces effectively and reduce the reliance on expensive style references. Experimental results demonstrate that DreamTalk is capable of generating photo-realistic talking faces with diverse speaking styles and achieving accurate lip motions, surpassing existing state-of-the-art counterparts.
1 Introduction
Audio-driven talking head generation, which concerns animating portraits with speech audio, has garnered significant interest due to its diverse applications in video games, film dubbing, and virtual avatars. Generating lifelike facial expressions is essential for enhancing the realism of talking heads [19]. These expressions during speech are termed as speaking styles [12, 46]. GANs [21] currently hold the state-of-the-art in expressive talking head generation [37, 30, 46, 20].However, their inherent issues with mode collapse and unstable training hamper their efficacy in consistently achieving high performance across a diverse range of speaking styles. Another issue is that prior methods often rely on reference videos [37, 30, 78] or texts [45, 90] to specify speaking styles. Their acquisition requires extra manual effort and hence is inflexible.
Figure 1: Leveraging the powerful diffusion models, DreamTalk is able to generate highly expressive talking heads across diverse speaking styles. Furthermore, DreamTalk is able to derive personalized speaking style directly from input audio, which obviates the need for additional style references.
As a new line of generative technique, diffusion models [26, 62] have recently been shown to produce high-quality results in numerous generative areas such as image generation [16, 53], video generation [60, 86, 27], and human motion synthesis [1, 72]. The success of diffusion models, stemming from their superior properties such as powerful distribution learning [16, 72], good convergence, and stylistic diversity, make them exceptionally promising for exploring expressive talking head generation. However, current diffusion-based talking head approaches [58, 66, 96] primarily concentrate on generating talking heads with neutral expressions and still struggle to produce satisfactory performance, e.g., suffering from frame jittering problem [58]. As a result, how to stimulate the full potential of diffusion models for expressive talking head generation is a promising yet untapped research direction.
In this paper, we propose DreamTalk, an expressive talking head generation framework that takes advantage of diffusion models to simultaneously deliver high performance across diverse speaking styles and reduce the reliance on expensive style references. Specifically, DreamTalk is composed of a denoising network, a style-aware lip expert, and a style predictor. The diffusion-based denoising network produces audio-driven facial motions with the speaking style specified by a reference video. The great distribution-learning characteristic of diffusion models endows the denoising network with the potential to produce high-quality results across diverse speaking styles. To harness this potential, we design a style-aware lip expert that drives the denoising network to produce accurate lip motions with vivid expressions. Contrasting with previous lip experts that overlook expression information and thus compromise style expressiveness, the proposed lip expert not only enhances lip accuracy but also ensures expressiveness. Finally, to further eliminate the need for additional style references, a diffusion-based style predictor is incorporated to predict personalized speaking styles directly from audio. The style predictor also incorporates the portrait as input during prediction, leveraging the correlation between speaker identity and speaking styles, thereby enhancing performance.
Ultimately, DreamTalk can consistently generate photo-realistic talking faces with precise lip-sync across a wide range of speaking styles while minimizing the need for additional style references. It also enables versatile manipulation of speaking styles and exhibits robust generalization across varied inputs, including songs, speech in multiple languages, noisy audio, and out-of-domain portraits. The effectiveness of DreamTalk is demonstrated through comprehensive qualitative and quantitative evaluations, showcasing its superiority over existing state-of-the-art methods.
2 Related Work
Audio-Driven Talking Head Generation. Audio-driven methods [14, 22, 81, 102, 67, 80] fall into two main categories: person-specific and person-agnostic. Person-specific approaches [68, 18, 29, 82, 41] are constrained to generating videos for speakers seen during training. Many of these [94, 74, 64, 35, 34, 29, 97, 98, 70] first craft 3D facial animations, later converting them into realistic videos. Recent advancements [23, 39, 57, 71, 93, 71] have employed neural radiance fields for modeling, yielding high-fidelity, realistic videos. Conversely, person-agnostic methods [10, 84, 55, 77] target generating videos for unseen speakers. Early methods prioritized lip synchronization [65, 5, 103, 6, 77, 49]. Later works shifted focus to natural facial expressions [95, 105] and head poses [7, 101, 83, 104, 100].
Expressive Talking Head Generation. Early methods [90, 20, 13, 24, 69, 82, 29, 61] model expressions in discrete emotion classes. To model more fine-grained expressions, most recent methods [30, 37, 46] leverage an expression reference video and transfer the expressions from that video to the generated one. However, these GAN-based models suffer from mode collapse, leading to videos with inferior lip-sync and style expressiveness. Our work addresses these issues by using diffusion models.
Specifying desired speaking styles effortlessly is also important for users. Most previous methods specify speaking styles using reference videos [30, 37, 46] or text [45, 90, 20], which needs human labor. A more user-friendly approach is to derive speaking styles from the input audio. Previous methods can only infer a limited number of discrete emotion classes from audio signals [29, 90, 61]. TH-PAD [95] generates expressions only aligned with the audio rhythm, not aligning with the emotional content of the audio. Besides, previous methods neglect the information in the input portrait. In this work, we aim to infer personalized and emotional expressions using input audio and portraits.
Diffusion Models. Diffusion models [62, 26] have demonstrated strong performance across multiple vision tasks [16, 53, 86, 99], including text-to-image generation [54, 31], image inpainting [42, 33], human motion generation [1, 72], 3D content generation [43], and video generation [60, 50, 89, 87, 79]. Previous efforts [58, 66, 3, 47] employing diffusion models in talking head generation only produce talking heads with neutral emotion and the results are unsatisfactory. Some methods devise diffusion-based renderers [91, 17] or face motion prior [96], yet still use GAN or regression models to model the audio-motion mapping. In this work, we endeavor to harness diffusion models for the generation of expressive talking heads, presenting a more intricate challenge with greater practical relevance.
3 Method
Figure 2: Illustration of DreamTalk. A style-aware lip expert (b), which evaluates the lip-audio synchronous probability under a given speaking style, is first trained to provide lip motion guidance for the denoising network (a). The denoising network is then trained to take the audio, the style reference video, and noisy face motion as input and predict the unnoised face motion. Then, A style predictor (c) is trained to predict the style code extracted from a video, taking audio and the speaker in that video as input. During inference process (d), the speaking style can be specified using style codes that are extracted from videos or derived from audio.
3.1 Problem Formulation
Given a portrait $\bm{I}$, a speech $\bm{A}$, and a style reference video $\bm{R}$, our method aims to generate a talking head video with lip motions synchronized with the speech and the speaking style reflected in the reference video. The audio $\bm{A}=[\bm{a}_{i}]_{i=1}^{L}$ is parameterized as a sequence of acoustic features. $\bm{R}$ is a sequence of video frames.
Besides, to eliminate the need for extra style references, our method also aims to infer the speaking style using solely the speech and the portrait. The inferred speaking style can replace the role of style reference videos in controlling the expressions, which enables our method to generate expressive talking head videos with only speech and portrait input.
3.2 DreamTalk
DreamTalk comprises 3 key components: a denoising network, a style-aware lip expert, and a style predictor.
The denoising network computes face motion conditioned on the speech and style reference video. The face motion $\bm{M}=[\bm{m}_{l}]_{l=1}^{L}$ is parameterized as a sequence of expression parameters from 3D Morphable Models [4]. The face motion is rendered into video frames by a renderer [52]. The style-aware lip expert provides lip motion guidance under diverse expressions and thus drives the denoising network to achieve accurate lip-sync while ensuring style expressiveness. The style predictor can predict the speaking style aligned with that conveyed in speech.
Denoising Network. The denoising network synthesizes face motion sequence frame-by-frame in a sliding window manner. It predicts a motion frame $\bm{m}_{l}$ using an audio window $\bm{A}_{w}=[\bm{a}_{i}]_{i=l-w}^{l+w}$, where $w$ denotes the window size.
The denoising network leverages forward and reverse diffusion processes. The diffusion process is modeled as a Markov noising process. Starting from a motion frame $\bm{m}_{(0)}$, it incrementally introduces Gaussian noise into the real data, gradually diffusing towards a distribution resembling $\mathcal{N}(\bm{0},\bm{I})$. Consequently, the distribution evolves as follows:
$q(\bm{m}_{(t)}|\bm{m}_{(t-1)})=\mathcal{N}(\sqrt{\alpha_{n}}\bm{m}_{(t-1)},(1-\alpha_{n})\bm{I}),$ (1)
where $\bm{m}_{(t)}$ is the motion frame sampled at diffusion step $t$, $t\in\{1,\dots,T\}$, and $\alpha_{n}$ is determined by the variance schedules. Conversely, the reverse diffusion process, or the denoising process, predicts the added noise in a noisy motion frame. Starting from a random motion frame $\bm{m}_{(T)}\sim\mathcal{N}(\bm{0},\bm{I})$, the denoising process incrementally removes the noise and recovers the original motion $\bm{m}_{(0)}$.
Instead of predicting the noise as formulated by [26], we follow [51] and predict the signal itself. The denoising network ${E}_{\theta}$ predicts $\bm{m}_{(0)}$ based on the noisy motion, the diffusion step, the speech context, and the style reference:
$\bm{m}^{*}_{(0)}=E_{\theta}(\bm{m}_{(t)},t,\bm{A}_{w},\bm{R}).$ (2)
The asterisk(***) indicates quantities that are generated.
Our denoising network has a transformer architecture [76]. The audio window $\bm{A}_{w}$ is first fed into a transformer-based audio encoder and the output is concatenated with the noisy motion $\bm{m}_{(t)}$ in the channel dimension. After linearly projected to the same dimension, the concatenated results and the timestep $t$ are summed and served as the key and value of a transformer decoder. To extract the speaking style from the style reference, a style encoder first extracts the sequence of 3DMM expression parameters from $\bm{R}$ and then feeds them into a transformer encoder. The output tokens are aggregated using a self-attention pooling layer [56] to obtain the style code $\bm{s}$. The style code is repeated $12w+$ 1 times and added with positional encodings. The results serve as the query of the transformer decoder. The middle output token of the decoder is fed into a feed-forward network to predict the signal $\bm{m}_{(0)}$.
Style-aware Lip Expert. We observe that using solely the denoising loss in standard diffusion models results in inaccurate lip motions. We conjecture that the loss alone is insufficient for the denoising network to effectively focus on generating precise lip motions. A typical remedy is to involve a pre-trained lip expert [49] that provides lip motion guidance. However, we observe the lip expert reduces the intensity of expressions. This stems from the fact that the lip expert merely focuses on a generic speaking style, which leads to generating face motions in a uniform style.
To address this issue, we introduce a style-aware lip expert. The proposed lip expert is trained to evaluate lip-sync under diverse speaking styles. Therefore, it can provide lip motion guidance under diverse speaking styles and strike a better balance between style expressiveness and lip-sync. The lip expert $\mathcal{E}$ computes the probability that a clip of audio and lip motions are synchronous conditioned on style reference $\bm{R}$:
$P_{{\text{sync}}}=\mathcal{E}([\bm{a}_{i}]_{i=l}^{l+n},[\bm{m}_{i}]_{i=l}^{l+n},\bm{R}),$ (3)
where $n$ denotes the clip length.
The style-aware lip expert encodes the lip motions and audio into respective embeddings conditioned on style reference and then computes the cosine similarity to represent the sync probability. To obtain lip motion information from face motion $\bm{m}$, we first convert $\bm{m}$ into the corresponding face mesh and select vertices in the mouth area as the representation of the lip motion [46]. The lip motion and audio encoders are mainly implemented by MLPs and 1D-convolutions, respectively. The style condition is fused into embeddings by first extracting style features from style reference using a style encoder, which mirrors the architecture of the one in the denoising network, and then concatenating the style features with intermediate feature maps from embedding encoders. The style encoder in the lip expert and the generator do not share parameters.
Style Predictor. Specifically, the style predictor $S_{\phi}$ predicts the style code $\bm{s}$ extracted by the style encoder in the trained denoising network. Observing the correlation between speaker identity and style codes (Sec. 4.4), the style predictor also integrates the portrait as input. The style predictor is instantiated as a diffusion model and is trained to predict the style code itself:
$\bm{s}^{*}_{(0)}=S_{\phi}(\bm{s}_{(t)},t,\bm{A},\bm{I}),$ (4)
where $\bm{s}_{(t)}$ is the style code sampled at diffusion step $t$.
The style predictor $S_{\phi}$ is a transformer encoder on a sequence consisting of, in order: audio embeddings, an embedding for the diffusion timestep, a speaker info embedding, the noised style code embedding, and a final embedding called learned query whose output is used to predict the unnoised style code. Audio embeddings are audio features extracted using self-supervised pre-trained speech models. To obtain the speaker info embedding, our method first extracts the 3DMM identity parameters, which include the face shape information but removes irrelevant information, such as expressions, from the portrait, and then embeds it into a token using an MLP.
Methods MEAD / HDTF / Voxceleb2
[…]
MakeItTalk [105] 0.73 / 0.57 / 0.52 0.11 / 0.24 / 0.24 3.97 / 5.12 / 6.29 5.32 / 4.55 / 5.15 2.10 / 3.16 / 2.17
Wav2Lip [49] 0.80 / 0.63 / 0.54 0.18 / 0.30 / 0.30 2.72 / 4.53 / 5.85 4.05 / 3.60 / 4.64 5.26 / 5.83 / 5.70
PC-AVS [104] 0.50 / 0.42 / 0.36 0.07 / 0.13 / 0.09 5.83 / 9.71 / 12.9 4.97 / 4.17 / 7.42 2.18 / 4.85 / 4.73
AVCT [84] 0.83 / 0.76 / 0.64 0.14 / 0.22 / 0.23 2.92 / 2.86 / 3.62 5.52 / 3.57 / 3.71 2.53 / 4.27 / 3.89
GC-AVT [37] 0.34 / 0.36 /    - 0.14 / 0.28 /    - 8.04 / 10.2 /    - 7.10 / 6.23 /    - 2.42 / 4.72 /    -
EAMM [30] 0.40 / 0.40 / 0.43 0.08 / 0.14 / 0.20 6.70 / 7.03 / 6.36 6.48 / 6.86 / 4.89 1.41 / 2.54 / 2.24
StyleTalk [46] 0.84 / 0.81 / 0.66 0.16 / 0.30 / 0.29 2.12 / 1.96 / 2.92 3.25 / 2.41 / 2.96 3.47 / 4.82 / 4.51
SadTalker [100] 0.69 / 0.77 / 0.44 0.16 / 0.24 / 0.19 4.12 / 5.99 / 9.12 4.37 / 4.07 / 6.11 2.76 / 4.35 / 4.38
PD-FGC [78] 0.49 / 0.41 / 0.35 0.05 / 0.13 / 0.12 5.50 / 9.50 / 12.5 4.10 / 4.23 / 8.19 2.27 / 4.68 / 4.64
EAT [20] 0.53 / 0.59 / 0.47 0.15 / 0.26 / 0.20 5.54 / 3.86 / 5.53 4.79 / 4.03 / 5.88 2.16 / 4.54 / 4.35
DreamTalk 0.86 / 0.85 / 0.69 0.16 / 0.31 / 0.30 1.93 / 1.80 / 2.69 2.91 / 2.15 / 2.72 3.78 / 5.17 / 4.90
Ground Truth 1 / 1 / 1 0.22 / 0.31 / 0.33 0 / 0 / 0 0 / 0 / 0 4.13 / 5.44 / 5.23
Table 1: Quantitative comparisons on MEAD, HDTF, and Voxceleb2. Since we only receive GC-AVT samples on MEAD and HDTF, GC-AVT is not evaluated on Voxceleb2.
3.3 Training and Inference
Training. The style-aware lip expert is first pre-trained by determining whether randomly sampled audio and lip motion clips are synchronous as in [49] and then frozen during training the denoising network.
The denoising network $E_{\theta}$ is trained by sampling random tuples $(\bm{m}_{(0)},t,\bm{A}_{w},\bm{R})$ from dataset, corrupting $\bm{m}_{(0)}$ into $\bm{m}_{(t)}$ by adding Gaussian noises, executing denoising steps to $\bm{m}_{(t)}$, and optimizing the loss:
$\mathcal{L}_{{\text{net}}}=\lambda_{{\text{denoise}}}\mathcal{L}_{{\text{denoise}}}+\lambda_{{\text{sync}}}\mathcal{L}_{{\text{sync}}}.$ (5)
Specifically, the ground-truth motion $\bm{m}_{(0)}$, and the speech audio window $\bm{A}_{w}$ are extracted from the training video of the same moment. $t$ is drawn from the uniform distribution $\mathcal{U}\{1,T\}$. The style reference $\bm{R}$ is a video clip randomly drawn from the same video containing $\bm{m}_{(0)}$.
We first compute the denoising loss of the diffusion models [26] defined as:
$\mathcal{L}_{{\text{denoise}}}=\lVert\bm{m}_{(0)}-E_{\theta}(\bm{m}_{(t)},t,\bm{A}_{w},\bm{R})\rVert_{2}^{2}.$ (6)
Then, the denoising network maximizes the synchronous probability via a sync loss on generated clips:
$\mathcal{L}_{{\text{sync}}}=-{\text{log}}(P_{{\text{sync}}}).$ (7)
Classifier-free guidance [25] is used to train our model. Specifically, $E_{\theta}$ is trained to learn both the style-conditional and unconditional distributions via randomly setting $\bm{R}=\varnothing$ by $10\%$ chance during training. $\varnothing$ is implemented as a sequence of face motions $[\bm{m}_{i}]$ with all zero values. For inference, the predicted signal is computed by
$\bm{m}^{*}_{(0)}$ $=\omega E_{\theta}(\bm{m}_{(t)},t,\bm{A}_{w},\bm{R})$ (8)
$+(1-\omega)E_{\theta}(\bm{m}_{(t)},t,\bm{A}_{w},\varnothing),$
instead of Equation 2. This approach enables controlling the effectiveness of the style reference $\bm{R}$ through adjustment of the scale factor $\omega$.
When training the style predictor, we draw a random video, then extract audio $\bm{A}$ and style code $\bm{s}_{(0)}$ (using the trained style encoder) from it. Since 3DMM identity parameters may leak expression information, the portrait $\bm{I}$ is sampled from another video with the same speaker identity. The style predictor $E_{\phi}$ is trained by optimizing the loss:
$\mathcal{L}_{{\text{pred}}}=\lVert\bm{s}_{(0)}-S_{\phi}(\bm{s}_{(t)},t,\bm{A},\bm{I})\rVert_{2}^{2},$ (9)
We utilize PIRenderer [52] as the renderer and meticulously fine-tune it to empower the renderer with emotional expression generation capabilities.
Inference. Our method enables the specification of speaking styles using either reference videos or solely through input audio and portrait. In the case of reference videos, style codes are derived using the style encoder in the denoising network. When relying solely on input audio and portrait, these inputs are processed by the style predictor, which employs a denoising procedure to obtain the style code.
With the style code, the denoising network utilizes the sampling algorithm of DDPM [26] to produce face motions. It first samples a random motion $\bm{m}_{(T)}^{*}\sim{}\mathcal{N}(\bm{0},\bm{I})$ then computes denoised sequences $\{\bm{m}_{(t)}^{*}\},{t=T-1,\dots,0}$ by incrementally removing the noise from $\bm{m}_{(t)}^{*}$. Finally, the motion $\bm{m}_{(0)}^{*}$ is the generated face motion. The sampling process can be accelerated by leveraging DDIM [63]. The output face motions are then rendered into videos by the renderer PIRenderer.
4 Experiments
4.1 Experimental Setup
Datasets. We train and evaluate the denoising network on MEAD [82], HDTF [101], and Voxceleb2 [11]. Since Voxceleb2 official videos are of low resolution, we redownload the original YouTube videos and re-crop the videos. The style-aware lip expert is trained on MEAD and HDTF. We train the style predictor on MEAD and evaluate it on MEAD and RAVEDESS [40].
Baselines. We compare our method with previous methods including: MakeitTalk [105], Wav2Lip [49], PC-AVS [104], AVCT [84], GC-AVT [37], EAMM [30], StyleTalk [46], DiffTalk [58], SadTalker [100], PD-FGC [78], and EAT [20]. For DiffTalk, since the released model is incomplete and unable to generate reasonable results until submission, we perform qualitative comparisons using videos from its released demo. For other methods, we generate the samples using released models or with the help of the authors.
Metrics. We utilize widely used metrics: SSIM [88], the Cumulative Probability of Blur Detection (CPBD) [48], the SyncNet confidence score ($\text{Sync}_{\text{conf}}$[9], the Landmark Distance around mouth area (M-LMD) [6], the Landmark Distance on the full face (F-LMD).
Figure 3: Qualitative comparisons with previous methods.
4.2 Main Results
Quanitative Comparisons. As shown in Tab. 1, our method outperforms previous methods across most metrics. Wav2Lip’s training with SyncNet as a discriminator explains its high SyncNet confidence score, even surpassing the ground truth. Notably, our method’s SyncNet confidence score closely aligns with the ground truth, and it achieves the best M-LMD scores, which indicates its capability for precise lip synchronization. Furthermore, our superior performance in the F-LMD metric demonstrates our method’s proficiency in generating facial expressions consistent with the reference speaking style.
Qualitative comparisons. Fig. 3 shows the qualitative comparisons. The portraits, style references, and audio are all unseen during training. It can be seen that MakeItTalk and AVCT struggle with accurate lip synchronization. While Wav2Lip and PC-AVS synchronize lips accurately, their outputs appear blurry. SadTalker, on the other hand, generally aligns lip movements with audio but occasionally displays unnatural jitters.
EAT’s capability is limited to generating discrete emotions, lacking the finesse for nuanced expressions. For example, in the left case, the style reference shows the speaker narrowing his eyes, but EAT merely produces a generic disgusted look with wide-open glaring eyes. Additionally, as shown in the right case, EAT struggles to maintain a consistent face shape during speaker head movements.
EAMM, GC-AVT, StyleTalk, and PD-FGC demonstrate the ability to produce fine-grained expressions. However, EAMM falls short in lip synchronization, GC-AVT and PD-FGC struggle with preserving speaker identity, and all three have issues rendering a plausible background. We observed that StyleTalk, while capable of generating nuanced expressions, occasionally does so with diminished intensity and fails to generate accurate lip motion for some words. A notable example is shown in the third column of the left case: when the speaker utters "um"; the expected closed-mouth motion is replaced by an open mouth in StyleTalk’s output. Besides, as shown in Fig. 4, StyleTalk occasionally generates videos where the face distorts suddenly.
Fig. 5 present comparisons with DiffTalk. DiffTalk struggles with lip synchronization and introduces jitteriness and artifacts in the mouth region.
In contrast, DreamTalk excels in producing realistic talking faces that not only mirror the reference speaking style but also achieve precise lip synchronization and superior video quality.
Figure 4: Face distortion observed in StyleTalk’s output.
Figure 5: Comparisons with DiffTalk.
Figure 6: The results generated using out-of-domain portraits.
Generalization Capabilities. Supplementary Video demonstrates our method’s capability to produce realistic talking head videos for out-of-domain portraits (paintings, toonified portraits, and sculptures, shown in Fig. 6), speech in various languages, noisy audio input, and songs. Sec. B.1 provides more analysis.
Figure 7: The results of speaking style prediction. The fourth column displays samples generated with predicted styles applied to the same portrait for clearer comparisons.
Results of Speaking Style Prediction. Fig. 7 presents the results of speaking style predictions. The speakers with angry emotion are sampled from MEAD while those with fear emotion are sampled from RAVEDESS. The style predictor, utilizing emotional audio and neutral portraits from test videos, adeptly deduces personalized speaking styles that align with those observed in the original videos. It demonstrates the capacity to discern subtle expressions within the same emotion. For instance, for samples with angry emotion, the first-row speaker exhibits narrowed eyes, in contrast to the second-row speaker’s intense, glaring stare. For samples with angry emotion, the first-row speaker’s eyes and mouth are open, whereas the second-row speaker combines narrowed eyes with a contorted facial expression.
The ablation study and user study for the style predictor are presented in Appendix A.
4.3 Ablation Study
Figure 8: Ablation study results
[…]
w/o Lip Expert 0.85 1.90 3.07 2.63
Uncond Lip Expert 0.83 2.19 3.42 4.51
Full 0.86 1.93 2.91 3.78
Table 2: The results of DreamTalk’s ablation study on MEAD. We omit CPBD scores since there are no significant differences between variants on CPBD.
To analyze the contributions of our designs, we conduct an ablation study with two variants: (1) remove the style-aware lip expert (w/o lip expert); (2) trained with unconditional lip expert (uncond lip expert). Our full model is denoted as Full.
Fig. 8 and Tab. 2 present our ablation study results. The variant w/o lip expert exhibits a decline in lip-sync accuracy on the emotional dataset MEAD, despite its competitive F-LMD score indicating expressive facial generation. Conversely, uncond lip expert secures a superior SyncNet confidence score at the expense of speaking style expressiveness. The Full model achieves a harmonious balance, ensuring both precise lip synchronization and vivid expressions, thanks to the style-aware lip expert directing the diffusion model’s expressive potential.
4.4 Style Code Visualization
Figure 9: t-SNE visualization of style codes from 15 speakers. Each color stands for style codes from an identical speaker.
Using t-distributed stochastic neighbor embedding (t-SNE) [75], we map style codes from the MEAD dataset’s 15 speakers into a 2D space. These speakers exhibit 22 distinct speaking styles, comprising seven emotions at three intensity levels, alongside a neutral style. For each style, style codes are extracted from 10 randomly chosen videos.
Fig. 9 reveals that style codes from identical speakers tend to cluster, despite some codes sharing identical emotion categories. This suggests that the variance in speaking styles attributable to individual speaker idiosyncrasies outweighs that induced by different emotions, thereby underpinning the rationale for using portrait information to infer speaking styles. We also observe that each speaker’s style code distribution exhibits both common patterns and individualized characteristics. We present these intriguing observations in Sec. B.2.
4.5 Style Manipulation
Figure 10: Effect of the input style controlled by adjusting the scale $\omega$ of the classifier-free guidance.
Adjusting the scale of Classifer-free Guidance. As elaborated in Sec. 3.3, the scale factor $\omega$ in the classifier-free guidance scheme modulates the effect of the input style. Adjusting $\omega$ either amplifies or attenuates the designated style, as shown in Fig. 10. When $\omega=0$, DreamTalk produces a talking head with a neutral expression. We observed that when the scale factor $\omega$ exceeds $2$, there is a noticeable decline in lip-sync accuracy.
Figure 11: The results of speaking style interpolation.
Style Code Interpolation. Leveraging the style space, we can modify speaking styles via style code manipulation. Fig. 11 illustrates that linear interpolation between style codes results in a seamless transition of generated speaking styles. This interpolation process allows for style intensity modulation and the generation of novel speaking styles.
4.6 User Study
[…]
MakeItTalk [105] 1.94 2.03 1.65
Wav2Lip [49] 2.29 1.45 1.14
PC-AVS [104] 2.26 1.81 1.86
AVCT [84] 2.31 2.21 1.72
GC-AVT [37] 2.35 1.20 1.88
EAMM [30] 1.81 1.40 1.78
StyleTalk [46] 2.35 2.29 2.08
SadTalker [100] 2.37 2.38 1.73
PD-FGC [78] 1.95 1.61 2.31
EAT [20] 2.24 1.65 2.29
DreamTalk 2.55 2.60 2.46
Ground Truth 3.03 2.89 1.83
Table 3: User study results.
We conduct a user study of 20 participants. We generate the test samples covering multiple speaking styles and speakers. For each method, the participant is required to score $10$ videos sampled from the test samples and is asked to give a rating (from 1 to 5, 5 is the best) on three aspects: (1) the Lip sync quality, (2) the realness of results, and (3) the style consistency between the generated videos and the style reference. As shown in Tab. 3, our method outperforms existing approaches across all aspects, particularly in style consistency, highlighting its superior capabilities.
5 Conclusion
In this work, we propose DreamTalk, a novel approach leveraging diffusion models for generating expressive talking heads. Our method aims to excel in diverse speaking styles while minimizing dependence on extra style references. We develop a denoising network for creating expressive, audio-driven facial motions and introduce a style-aware lip expert to optimize lip-sync without compromising style expressiveness. Additionally, we devise a style predictor that infers speaking styles directly from audio, eliminating the need for video references. The efficacy of DreamTalk is validated through extensive experiments.
Acknowledgements. This work is supported by Alibaba Group through Alibaba Research Intern Program. We would like to thank Xinya Ji, Borong Liang, Yan Pan, and Suzhen Wang for their generous help with the comparisons.
Supplementary Material
Appendix A Additional Evaluation for Speaking Style Prediction
A.1 Ablation Study
Figure 12: The qualitative results of style predictor’s ablation study.
To evaluate the impact of our design choices, we conduct an ablation study with three variants: (1) omitting speaker information and relying solely on audio for prediction (w/o speaker info); (2) during model training, the speaker info and audio are both obtained from the same video (w/o cross-ID training); (3) employing a regression model instead of a diffusion model for prediction (regression). Our full model is denoted as Full. When generating samples for evaluation, the facial images and audio we use are sourced from videos of the same individual expressing different emotions(e.g. the face image is from a happy video while the audio is from an angry one.). This generation approach better aligns with real-world applications.
How to quantitatively evaluate the performance of speaking style prediction has not been explored before. we devise three metrics:
  • Style Code Distance (SCD) We extract the style codes from the videos that provide the audio input and compute the L2 distance between the predicted style codes and these style codes.
  • Motion Distance (MD) We use the predicted style codes and the audio used for prediction to generate face motions and compute the L2 distance between the generated face motions and the face motions extracted from the ground truth videos.
  • Style Accuracy (SA) We split test videos into different speaking styles and train a style classifier to classify which style a face motion sequence belongs to. Then, We classify the face motions generated using predicted style codes and report the accuracy. Specifically, we put the videos from the same speaker, emotion, and intensity into one style. We evaluate this metric on MEAD only since the number of RAVEDESS videos for each style is inadequate to train a style classifier. The ground truth testing set gets $92.5\%$ accuracy.
We refrain from devising image-level metrics, such as training an image classifier for speaking style classification, due to several critical considerations. Firstly, factors in images that are irrelevant to expression, such as the speaker’s identity and background elements, can adversely impact the accurate prediction of nuanced speaking styles. Secondly, inaccuracies introduced by the rendering process may further additionally hinder the accurate discernment of these subtle speaking styles.
[…]
w/o speaker info 0.49 0.28 64.3
w/o cross-ID training 0.68 0.45 28.1
regression 0.56 0.32 55.1
Full 0.42 0.23 78.6
Table 4: The ablation study results of the style predictor.
The results are shown in Tab. 4 and Fig. 12. The w/o speaker info variant successfully predicts emotions from audio but occasionally fails to maintain consistency between the predicted speaking style and speaker identity, leading to poor identity preservation. This underscores the importance of speaker information in predicting speaking styles. Although in experiments, we observed that w/o cross-ID training achieves slightly better performance than Full when the input portrait and audio are from the same video, it underperformed, often failing to predict the correct emotion, when inputs were from different videos. This suggests that identity 3DMM parameters may convey some expression information, and without cross-ID training, the model might derive emotional cues from this leaked information rather than the audio. The regression variant struggles to generate accurate expressions for certain data, highlighting the superior distribution-learning capability of diffusion models in facilitating speaking style prediction.
A.2 User Study
In our user study, we evaluate the alignment between the original and predicted speaking styles. Directly assessing the alignment of speaking styles can be somewhat ambiguous, so we employ a comparative approach for evaluation. Specifically, we create a series of video triplets. Each triplet consisted of a test video from our dataset and two generated videos. The first video was generated using a style code predicted from an input portrait, sharing the same speaker identity as in the test video but displaying a neutral emotion, combined with the audio from the test video. The second video is generated using the style code extracted from videos with the same emotion but from a speaker different from the one in the test video. We recruit 20 participants. Each participant is then asked to evaluate 20 triplets and identify which of the generated videos most accurately reflected the speaking style of the test video. The videos generated using predicted style codes are preferred in $75.8\%$ of all ratings. This indicates that the style predictor is able to infer personalized speaking styles that are aligned with the audio.
A.3 Analyzing the Influence of Portraits
Figure 13: Analyzing the influence of portraits on style prediction. The audio conveys surprised emotion.
We analyze the influence of portraits on speaking style prediction by predicting speaking styles with an audio clip and different input portraits. The predicted styles are subsequently applied to an identical portrait for a clearer comparison. As shown in Fig. 13, the predicted speaking styles match the subtle identity characteristics, such as gender, of the input portraits. The predicted style A generated more feminine results. This validates the necessity of integrating portrait information during style prediction.
Appendix B Additional Results for Expressive Talking Head Generation
B.1 Analysis on Generalization Capabilities
Songs. As demonstrated in Supplementary Video, our method successfully generates reasonable results for songs, even those with accompaniment, despite this being significantly different from the training dataset’s data distribution. A noticeable decline in lip-sync accuracy is observed when the accompaniment volume is excessively high. We conduct a comparative analysis of lip-sync performance between songs with accompaniment and songs with removed accompaniment (using data from the SingFace Dataset [38]). It is found that the accompaniment adversely affects lip-sync, leading to mouth movements resembling mumbling. Addressing the negative impact of accompaniment on lip-sync accuracy presents an interesting avenue for future research.
Speech in Multiple Languages. Supplementary Video shows that our method generates satisfactory results with speech in French, Chinese, Spanish, German, Italian, Japanese, and Korean. The versatility of wav2vec features aids application across various languages. Additionally, the inclusion of multilingual talking head videos from Voxceleb2 enhances generalization.
Noisy Audio. Supplementary Video demonstrates that our method yields satisfactory outcomes when processing audio mixed with multiple noise types and intensities. We employ noise recordings from typical talking head application environments—meetings, offices, and cafeterias—sourced from the DEMAND dataset [73]. Using a publicly available tool [^1] https://github.com/Sato-Kunihiko/audio-SNR, we blend the audio with noise at SNRs of 20 dB, 10 dB, and 0 dB. Remarkably, our method maintains performance even at 0 dB SNR, where the noise is as loud as the speech and significantly impairs speech intelligibility and clarity.
B.2 More Results of Style Code Visualization
Figure 14: t-SNE visualization of style codes for 3 speakers, with darker hues representing increased emotional intensity.
We observe that each speaker’s style code distribution exhibits both common patterns and individualized characteristics. Common patterns include: Firstly, speaking styles of different emotions cluster together first, with styles of lower intensity being closer to neutral and those of higher intensity being further away. Secondly, speaking styles of anger and disgust, as well as fear and surprise, often cluster together, as shown in Fig. 14 (b) and (c). Note that unlike Ma et al. [46], our method does not incorporate losses to constrain style space.
Fig. 14 (a) illustrates an example of individualized characteristics. The speaker’s manifestation of fear closely resembles sadness, lacking the characteristic wide-eyed and open-mouthed expression, thereby positioning the speaking styles of fear nearer to those of sadness rather than surprise. Even within the same emotion, a speaker’s speaking style can exhibit notable variation. In Fig. 14 (c), the speaker’s dual expression of sadness—once with clenched teeth, similar to happy expressions, and another with depressed lip corners, akin to disgust—results in style codes close to the respective emotions. This observation diminishes the rationale for manually categorizing styles based on emotion and intensities in Ma et al. [46].
Appendix C Implementation Details
C.1 Architectual Details
Denoising Network.The audio encoder processes an input window of 11 sequential audio features, each of dimension $1024$. These features undergo dimension reduction to $256$ via a linear layer and then are fed to a transformer encoder comprising three 8-head transformer encoder layers, each with a hidden size of 256. Subsequently, a linear layer transforms the output tokens to yield audio tokens sized $11\times 256$. The audio tokens are concatenated with noisy motion and then added with the encoded diffusion step.
The style encoder ingests sequential expression parameters from style reference videos, each sequence sized $N\times 64$. These sequences, ranging in length from 64 to 256 frames, are initially expanded to 256 dimensions via a linear layer. Subsequently, they are introduced into a transformer encoder, composed of three 8-head layers, each with a hidden size of 256. The resulting output tokens, each with a dimension of $256$, are aggregated through self-attention pooling [56], yielding a style code of dimension $256$.
Within the decoder, the style code is repeated $11$ times, subsequently added with positional embedding to produce style tokens. These tokens, in conjunction with audio tokens, are processed by a transformer decoder, encompassing three 8-head layers, each with a hidden dimension of 256. Here, style tokens serve as the query, while audio tokens serve as both key and value. The middle output token is fed into a linear layer to predict facial motion.
Style-aware Lip Expert.
Figure 15: The architecture of mouth embedder and audio embedder. ConvBlock consists of a 1D CNN layer and a batch normalization layer. "k3, 32, s2" means that the kernel size is 3, the output dimension is 32, and the stride is 2. ConvRBlock is a ConvBlock with residual connection.
The face mesh is obtained by adding the mean shape to the product of the expression parameters and expression bases. Fig. 15 shows the architecture of the audio embedder and the mouth embedder.
Style Predictor. The style predictor is implemented as a transformer encoder comprising six 8-head transformer encoder layers, each with a hidden size of $256$. The input features are all linearly projected to $256$.
C.2 Data Details
C.2.1 Datasets
MEAD. The dataset, an in-lab talking-face corpus, features 60 speakers articulating eight emotions at three different intensity levels. When dividing the MEAD dataset into training and test subsets, we adhere to previously established methodologies [30].
HDTF. The dataset stands out as a high-resolution, in-the-wild audio-visual dataset. We designate $10\%$ of HDTF videos for testing and reserve the remainder for training.
Voxceleb2. Voxceleb2 is a large-scale talking head dataset with videos collected from YouTube. We redownload and recrop the videos to improve their resolution to $256\times 256$. Subsequently, approximately $80000$ high-quality videos are selected, with $400$ allocated for testing and the rest for training.
RAVDESS. The dataset features 24 professional actors (12 female, 12 male) vocalizing two lexically matched statements in a neutral North American accent. It encompasses a range of expressions in both speech and song, each articulated at two levels of emotional intensity, with an additional neutral expression included. We employ the speech data from RAVDESS, encompassing eight emotions, for evaluation.
C.2.2 Data Processing
The original videos are cropped and resized to $256\times 256$ pixels, aligning with the specifications in FOMM [59], and are sampled at 25 FPS. The 3DMM parameters are extracted by Deep3DFace [15].
Regarding the audio features used in the denoising network and the style-aware lip expert, we downsample the speech wave into the sampling rate of 16000 and extract acoustic features employing a pre-trained Wav2Vec2.0 model [2].
For audio features used in the style predictor, we extract them using a pre-trained HuBERT model [28]. Besides, we also utilize low-level audio features including Mel Frequency Cepstrum Coefficients (MFCC), Mel-filterbank energy features (FBANK), fundamental frequency, and voice flag. These two type of features are concatenated to represent the audio features used in the style predictor.
C.3 Training Details
Our framework is implemented on Pytorch. We employ Adam [32] for optimization, with a learning rate set to $0.0001$. The number of diffusion steps for the denoising network and style predictor is $1000$. The training batch size for the denoising network, style predictor, and lip expert is $64$, $64$, and $32$, respectively. $\lambda_{\text{denoise}}$, $\lambda_{\text{sync}}$, $n$, and $w$ are set to $1$, $1$, $5$, and $5$, respectively. The number of frames for style reference and audio used in style prediction is limited to $64-256$, corresponding to a time length of $2.56-10.24$ seconds. The denoising network, style predictor, and lip expert are trained on one NVIDIA Tesla A100 GPU for about $3$, $1$, and $10.5$ hours, respectively.
C.3.1 Finetuning PIRender
The renderer is fine-tuned with the losses in  Ren et al. [52] using MEAD. Instead of training with the self-reconstruction protocol where the source frame and target frame are from the same video, we select the source frame and target frame from the same speaker with different emotions. This enables the renderer to generate emotions different from the input portrait. This also allows our method to utilize portraits with emotions, unlike previous approaches that are confined to using neutral portraits [30]. We also observed that when fine-tuned only on the emotional talking head dataset, the renderer struggles with identity preservation. We argue that this problem stems from the fact that the number of speakers in the emotional dataset is limited. Therefore, we incorporate some neutral videos in Voxceleb into the data used for fine-tuning. This enhances the performance in identity preservation.
C.3.2 Training Style-aware Lip Expert
The style-aware lip expert is trained to discriminate whether the input audio and face motions are synchronized. We use cosine-similarity with binary cross-entropy loss to train the lip expert. Specifically, we compute cosine-similarity for the face motion embedding $\bm{e}^{m}$ and audio embedding $\bm{e}^{a}$ to represent the probability that the input audio-motion pair is synchronized. The training loss of the lip expert is:
$\mathcal{L}_{{\text{expert}}}={\text{BCE}}(\frac{\bm{e}^{m}\cdot\bm{e}^{a}}{\max(||\bm{e}^{m}||_{2}\cdot||\bm{e}^{a}||_{2},\epsilon)}),$ (10)
where $\epsilon$ is a small number for avoiding the division-by-zero error.
C.4 Inference Details
The inference of the denoising network can be accelerated with DDIM. We generate samples with $10$ DDIM steps and observe no performance drop. Generating a 30-second video offline takes $15.61$ seconds, with the face motion generation only taking $1.24$ seconds. During evaluation, the scale factor $\omega$ of classifier-free guidance is set to $1$. The style predictor uses the sampling algorithm of DDPM to predict style codes.
The emotion conveyed in the style reference (video or audio), should remain consistent to avoid confusing the model.
The head pose information, which is fed into the renderer, can be derived from real videos or generated using existing methods [83].
Appendix D Limitations and Future Work
Despite DreamTalk’s promising advancements in expressive talking head generation, it encounters several challenges that open avenues for future research.
Firstly, the method occasionally produces artifacts, such as teeth flickering, around the mouth area, particularly during intense expressions. Generating teeth is a long-standing challenge in talking head generation since the algorithm needs to inpaint the teeth area that are often occluded in the input portrait. This problem is exacerbated under intense expressions where the teeth area expands. The issue can be mitigated by incorporating modules [44, 92, 36] proposed recently that enhance the teeth quality. A more comprehensive solution involves developing an emotion-specific renderer, as current renderers [90, 46, 20, 30] are mainly adaptations of existing facial reenactment methods with minimal modifications. An emotion-aware renderer would not only address the teeth generation issue but also enhance the overall expressiveness of emotions.
Secondly, DreamTalk does not account for the temporal variability in speaking styles. In real-life scenarios, a speaker’s style evolves over time, a feature our method currently overlooks. For instance, at the end of a speech, our method might still produce expressions of intense emotion, such as a wide-open mouth in surprise, instead of a more neutral, closed-mouth expression. Introducing a module that dynamically predicts speaking style over time could address this limitation.
Thirdly, the style predictor sometimes struggles with accurately identifying emotions in low-intensity audio clips from the MEAD dataset (some audio in MEAD intensity level 1 videos). The emotion conveyed in these audio clips is very similar to neutral emotion and thus confusing the prediction. We also observed that although the videos in the MEAD dataset clearly express emotions due to the good training and supervision of the speakers, sometimes the audio does not correspond with the emotions that should be expressed. Therefore, To enhance prediction accuracy, employing a dataset where the audio closely aligns with the expressed emotions could be beneficial. Another solution is to incorporate textual information from audio during prediction, a strategy commonly employed in speech emotion recognition [85, 8].
Despite these challenges, DreamTalk marks a significant stride in the realm of high-quality, expressive talking head generation, setting a foundation for further innovations.
Appendix E Ethical Consideration
DreamTalk is able to generate realistic talking head videos. This positions DreamTalk with a broad spectrum of potential applications, each carrying intricate societal implications. While DreamTalk holds significant potential in amplifying and enriching human creative endeavors and may pave the way for innovative tools for creative professionals, its capabilities also harbor risks. There’s a possibility for DreamTalk to generate content that might encompass or imply sexual themes, promote hatred, or depict violence. Misuse of DreamTalk could lead to negative repercussions on individuals or groups, potentially erasing or maligning them, perpetuating stereotypes, and subjecting them to disrespect. Other concerns include the potential for harassment, intimidation, or exploitation. Furthermore, DreamTalk’s capabilities might be harnessed to mislead or spread misinformation.
Before releasing DreamTalk, we have implemented and plan to introduce several safeguards to curb potential misuse. We’ve purged detrimental content from the training dataset and incorporated visual filters to deter users from creating harmful outputs. To counteract biases in the generated results, we’re enhancing the dataset’s diversity by manually ensuring balance, which will reduce instances of erasure, stereotype perpetuation, indignity, and uneven quality across inputs. Users will be advised against using images without the depicted individuals’ consent to combat harassment and bullying. To prevent the spread of misinformation, all DreamTalk outputs will bear watermarks indicating their synthetic nature. Our pre-release strategy involves a thorough risk assessment, leveraging a growing suite of safety evaluations and red teaming techniques. We’ll also scrutinize the findings from pilot tests centered on new use cases and conduct in-depth post-release assessments. Both automated and manual monitoring mechanisms are in development to preempt misuse. Our commitment remains steadfast in continuously researching ways to minimize adverse societal effects.

### v2

@@
-DreamTalk: When Expressive Talking Head Generation Meets Diffusion Probabilistic Models
+DreamTalk: When Emotional Talking Head Generation Meets Diffusion Probabilistic Models
 Abstract
-Diffusion models have shown remarkable success in a variety of downstream generative tasks, yet remain under-explored in the important and challenging expressive talking head generation.
-In this work, we propose a DreamTalk framework to fulfill this gap, which employs meticulous design to unlock the potential of diffusion models in generating expressive talking heads.
+Emotional talking head generation has attracted growing attention.
+Previous methods, which are mainly GAN-based, still struggle to consistently produce satisfactory results across diverse emotions and cannot conveniently specify personalized emotions.
+In this work, we leverage powerful diffusion models to address the issue and propose DreamTalk, a framework that employs meticulous design to unlock the potential of diffusion models in generating emotional talking heads.
 Specifically, DreamTalk consists of three crucial components: a denoising network, a style-aware lip expert, and a style predictor.
-The diffusion-based denoising network is able to consistently synthesize high-quality audio-driven face motions across diverse expressions.
-To enhance the expressiveness and accuracy of lip motions, we introduce a style-aware lip expert that can guide lip-sync while being mindful of the speaking styles.
-To eliminate the need for expression reference video or text, an extra diffusion-based style predictor is utilized to predict the target expression directly from the audio.
-By this means, DreamTalk can harness powerful diffusion models to generate expressive faces effectively and reduce the reliance on expensive style references.
-Experimental results demonstrate that DreamTalk is capable of generating photo-realistic talking faces with diverse speaking styles and achieving accurate lip motions, surpassing existing state-of-the-art counterparts.
-1 Introduction
-Audio-driven talking head generation, which concerns animating portraits with speech audio, has garnered significant interest due to its diverse applications in video games, film dubbing, and virtual avatars.
-Generating lifelike facial expressions is essential for enhancing the realism of talking heads [19].
-These expressions during speech are termed as speaking styles [12, 46].
-GANs [21] currently hold the state-of-the-art in expressive talking head generation [37, 30, 46, 20].However, their inherent issues with mode collapse and unstable training hamper their efficacy in consistently achieving high performance across a diverse range of speaking styles.
-Another issue is that prior methods often rely on reference videos [37, 30, 78] or texts [45, 90] to specify speaking styles.
-Their acquisition requires extra manual effort and hence is inflexible.
-Figure 1: Leveraging the powerful diffusion models, DreamTalk is able to generate highly expressive talking heads across diverse speaking styles.
-Furthermore, DreamTalk is able to derive personalized speaking style directly from input audio, which obviates the need for additional style references.
-As a new line of generative technique, diffusion models [26, 62] have recently been shown to produce high-quality results in numerous generative areas such as image generation [16, 53], video generation [60, 86, 27], and human motion synthesis [1, 72].
-The success of diffusion models, stemming from their superior properties such as powerful distribution learning [16, 72], good convergence, and stylistic diversity, make them exceptionally promising for exploring expressive talking head generation.
-However, current diffusion-based talking head approaches [58, 66, 96] primarily concentrate on generating talking heads with neutral expressions and still struggle to produce satisfactory performance, e.g., suffering from frame jittering problem [58].
-As a result, how to stimulate the full potential of diffusion models for expressive talking head generation is a promising yet untapped research direction.
-In this paper, we propose DreamTalk, an expressive talking head generation framework that takes advantage of diffusion models to simultaneously deliver high performance across diverse speaking styles and reduce the reliance on expensive style references.
+The diffusion-based denoising network can consistently synthesize high-quality audio-driven face motions across diverse emotions.
+To enhance lip-motion accuracy and emotional fullness, we introduce a style-aware lip expert that can guide lip-sync while preserving emotion intensity.
+To more conveniently specify personalized emotions, a diffusion-based style predictor is utilized to predict the personalized emotion directly from the audio, eliminating the need for extra emotion reference.
+By this means, DreamTalk can consistently generate vivid talking faces across diverse emotions and conveniently specify personalized emotions.
+Extensive experiments validate DreamTalk’s effectiveness and superiority.
+The code is available at https://github.com/ali-vilab/dreamtalk.
+Index Terms:
+Emotional talking head generation, Diffusion models
+publicationid: pubid:
+I Introduction
+Audio-driven talking head generation, which concerns animating portraits with speech audio, has garnered significant interest due to its diverse applications, such as film dubbing, digital human generation, video conferences in band-limited conditions, and online education.
+To produce realistic talking heads, it is crucial to generate and control emotions.
+Recognizing that a single type of emotion can be expressed in diverse ways, recent research focus has shifted from modeling coarse-grained, discrete emotions to modeling more fine-grained, personalized emotions [1, 2, 3].
+These personalized emotions are also called speaking styles [2, 4].
+A speaking style is defined as facial motion patterns reflected in a talking video clip.
+In different video clips, the speakers may have different speaking habits and may display various emotions.
+Therefore, speaking styles are diverse.
+Existing methods still struggle to 1) consistently produce high-quality results across diverse speaking styles and 2) conveniently specify desired speaking styles.
+Existing methods [3, 5, 2, 6] are mainly based on GANs [7].
+GANs’ inherent issues, such as mode collapse and unstable training, impair their performance across diverse speaking styles.
+Although these methods achieve satisfactory results for a limited range of speaking styles, they struggle with more diverse styles, especially ones unseen during training, often resulting in diminished emotional intensity, inaccurate lip motion, or sudden facial distortions [2].
+Another issue is that to specify speaking styles, previous methods often rely on extra references, such as videos [3, 5, 1] or texts [8, 9].
+Their acquisition requires extra manual effort and hence is inconvenient.
+Figure 1: Leveraging the powerful diffusion models, DreamTalk can consistently generate high-quality talking heads across diverse speaking styles.
+Furthermore, DreamTalk can conveniently use audio to specify personalized speaking style, obviating the need for additional style references.
+As a new line of generative technique, diffusion models [10, 11] have shown capability to produce high-quality results in numerous generative areas [12, 13, 14, 15, 16].
+The success of diffusion models, stemming from their superior properties such as powerful distribution learning [12, 17], make them exceptionally promising for exploring emotional talking head generation.
+However, current diffusion-based talking head approaches [18, 19, 20, 21] primarily concentrate on generating talking heads with neutral expressions or a limited number of discrete emotions, lacking diverse and fine-grained speaking styles.
+Therefore, exploring the full potential of diffusion models for generating talking heads with diverse speaking styles represents a promising, yet unexplored, research direction.
+In this paper, we propose DreamTalk, an emotional talking head generation framework that takes advantage of diffusion models to consistently deliver high performance across diverse speaking styles and reduce the reliance on expensive style references.
 Specifically, DreamTalk is composed of a denoising network, a style-aware lip expert, and a style predictor.
 The diffusion-based denoising network produces audio-driven facial motions with the speaking style specified by a reference video.
-The great distribution-learning characteristic of diffusion models endows the denoising network with the potential to produce high-quality results across diverse speaking styles.
-To harness this potential, we design a style-aware lip expert that drives the denoising network to produce accurate lip motions with vivid expressions.
-Contrasting with previous lip experts that overlook expression information and thus compromise style expressiveness, the proposed lip expert not only enhances lip accuracy but also ensures expressiveness.
-Finally, to further eliminate the need for additional style references, a diffusion-based style predictor is incorporated to predict personalized speaking styles directly from audio.
-The style predictor also incorporates the portrait as input during prediction, leveraging the correlation between speaker identity and speaking styles, thereby enhancing performance.
-Ultimately, DreamTalk can consistently generate photo-realistic talking faces with precise lip-sync across a wide range of speaking styles while minimizing the need for additional style references.
-It also enables versatile manipulation of speaking styles and exhibits robust generalization across varied inputs, including songs, speech in multiple languages, noisy audio, and out-of-domain portraits.
-The effectiveness of DreamTalk is demonstrated through comprehensive qualitative and quantitative evaluations, showcasing its superiority over existing state-of-the-art methods.
-2 Related Work
+The great distribution-learning property of diffusion models enable the denoising network to consistently produce high-quality results across diverse speaking styles.
+To enhance the lip-sync, we design a style-aware lip expert that drives the denoising network to produce accurate lip motions under different speaking styles.
+We observe that previous lip experts, which neglect emotional information, compromise the intensity of generated emotions.
+To preserve the emotion intensity, we find it important to integrate style information into the lip expert, thereby making it style-aware.
+Finally, to eliminate the need for additional style references, a diffusion-based style predictor is incorporated to predict personalized speaking styles directly from audio.
+To predict more personalized emotions, we find it crucial to leverage the correlation between speaker identity and speaking styles; therefore, we provide identity information by incorporating the portrait as input.
+The effectiveness of DreamTalk is demonstrated through comprehensive qualitative and quantitative evaluations.
+DreamTalk can even generate reasonable results for songs in multiple languages, despite these audios being significantly different from those in the training set.
+In summary, our contributions are as follows:
+  • We propose DreamTalk, a diffusion-based framework that can consistently generate talking faces with precise lip-sync as well as rich emotions across diverse speaking styles.
+We find that diffusion models achieve better results than GANs for more diverse speaking styles.
+  • We explore how to use audio alone to predict personalized emotions, making it more convenient than relying on extra videos to specify speaking styles.
+We discover that incorporating identity information significantly enhances prediction accuracy.
+  • We propose a style-aware lip expert that can avoid reducing emotion intensity when providing lip guidance.
+We find that making the lip expert conditioning on speaking style information is crucial for maintaining emotional fullness.
+  • Trained in a classifier-free manner, DreamTalk can use the classifier-free guidance scheme to adjust the intensity of arbitrary speaking styles.
+II Related Work
 Audio-Driven Talking Head Generation.
@@
 Later works shifted focus to natural facial expressions [19, 55] and head poses [56, 57, 58, 59, 60].
-Expressive Talking Head Generation.
-Early methods [90, 20, 13, 24, 69, 82, 29, 61] model expressions in discrete emotion classes.
-To model more fine-grained expressions, most recent methods [30, 37, 46] leverage an expression reference video and transfer the expressions from that video to the generated one.
-However, these GAN-based models suffer from mode collapse, leading to videos with inferior lip-sync and style expressiveness.
+FROND [61] introduces a fine-grained motion model that captures local facial movement keypoints and embeds overall motion context to predict audio-driven facial movements and achieve smooth temporal transitions.
+However, this method fails to generate emotional expressions during speech, thereby affecting the video’s realism.
+Emotional Talking Head Generation.
+Early methods [9, 6, 62, 63, 64, 31, 30, 65, 66] model expressions in discrete emotions.
+To model fine-grained emotions, recent methods [5, 3, 2, 67] use an expression reference video and transfer the expressions from that video to the generated one.
+However, these GAN-based methods cannot consistently achieve high performance across diverse emotions.
 Our work addresses these issues by using diffusion models.
-Specifying desired speaking styles effortlessly is also important for users.
-Most previous methods specify speaking styles using reference videos [30, 37, 46] or text [45, 90, 20], which needs human labor.
+UniFaceGAN [68] introduces a temporally consistent facial video editing framework that handles both face swapping and face reenactment simultaneously by using a 3D reconstruction model and a novel temporal loss constraint.
+Facial-Prior-Guided FME Generation [69] enhances facial micro-expression generation by utilizing adaptive weighted prior maps and facial priors to guide motion representation.
+F3A-GAN [70] employs a 3D geometric flow, termed facial flow, to represent natural facial motion for continuous image synthesis.
+Although all these methods are related to facial expression generation, none of them can generate accurate lip shapes driven by audio in different emotional contexts.
+Conveniently specifying desired speaking styles is also important.
+Most previous methods rely on reference videos [5, 3, 2] or text [8, 9, 6], which needs human labor.
 A more user-friendly approach is to derive speaking styles from the input audio.
-Previous methods can only infer a limited number of discrete emotion classes from audio signals [29, 90, 61].
+Previous methods can only infer a limited number of discrete emotion classes from audio  [30, 9, 65].
 TH-PAD [19] generates expressions only aligned with the audio rhythm, not aligning with the emotional content of the audio.
-Besides, previous methods neglect the information in the input portrait.
-In this work, we aim to infer personalized and emotional expressions using input audio and portraits.
+Besides, previous methods neglect information in the input portrait.
+In this work, we aim to infer personalized emotions using input audio and portraits.
 Diffusion Models.
-Diffusion models [62, 26] have demonstrated strong performance across multiple vision tasks [16, 53, 86, 99], including text-to-image generation [54, 31], image inpainting [42, 33], human motion generation [1, 72], 3D content generation [43], and video generation [60, 50, 89, 87, 79].
-Previous efforts [58, 66, 3, 47] employing diffusion models in talking head generation only produce talking heads with neutral emotion and the results are unsatisfactory.
-Some methods devise diffusion-based renderers [91, 17] or face motion prior [96], yet still use GAN or regression models to model the audio-motion mapping.
-In this work, we endeavor to harness diffusion models for the generation of expressive talking heads, presenting a more intricate challenge with greater practical relevance.
-3 Method
+Diffusion models [11, 10] have demonstrated strong performance across multiple vision tasks [12, 14, 71, 72, 73, 74], including text-to-image generation [75], human motion generation [17], and video generation [15, 76, 77, 78, 79].
+Most diffusion-based methods for talking head generation [18, 80, 81, 82, 83, 84, 19, 85, 86, 87, 88], including EMO [21], AniPortrait [85] and Hallo [89], mainly generate talking heads with neutral emotions and lacks emotional controllability.
+Besides, the inference of EMO is slow.
+VASA [20] can only generate a limited number of emotions, lacking diverse, fine-grained speaking styles.
+In this work, we aim to harness diffusion models for generating and controlling diverse, fine-grained speaking styles in talking heads, presenting a more intricate challenge.
+III Method
 Figure 2: Illustration of DreamTalk.
-A style-aware lip expert (b), which evaluates the lip-audio synchronous probability under a given speaking style, is first trained to provide lip motion guidance for the denoising network (a).
-The denoising network is then trained to take the audio, the style reference video, and noisy face motion as input and predict the unnoised face motion.
-Then, A style predictor (c) is trained to predict the style code extracted from a video, taking audio and the speaker in that video as input.
-During inference process (d), the speaking style can be specified using style codes that are extracted from videos or derived from audio.
-3.1 Problem Formulation
+A style-aware lip expert (b) is first trained to provide lip motion guidance for the denoising network (a).
+The denoising network is then trained to predict emotional audio-driven face motions.
+Then, A style predictor (c) is trained to use audio to predict the style code.
+During inference (d), the speaking style can be specified using style codes that are extracted from videos or derived from audio.
+III-A Problem Formulation
 Given a portrait $\bm{I}$, a speech $\bm{A}$, and a style reference video $\bm{R}$, our method aims to generate a talking head video with lip motions synchronized with the speech and the speaking style reflected in the reference video.
@@
 $\bm{R}$ is a sequence of video frames.
-Besides, to eliminate the need for extra style references, our method also aims to infer the speaking style using solely the speech and the portrait.
-The inferred speaking style can replace the role of style reference videos in controlling the expressions, which enables our method to generate expressive talking head videos with only speech and portrait input.
-3.2 DreamTalk
-DreamTalk comprises 3 key components: a denoising network, a style-aware lip expert, and a style predictor.
+The head motions in the generated videos can originate from real videos or be produced by existing methods [60, 4].
+Besides, to conveniently specify speaking styles, our method also aims to infer the speaking style using solely the speech and the portrait, obviating the need for extra style references.
+The inferred speaking style can replace the role of style reference videos in controlling expressions (fig.
+1), enabling our method to generate personalized emotions with only speech and portrait.
+III-B DreamTalk
+As illustrated in fig. 2, DreamTalk comprises 3 key components: a denoising network, a style-aware lip expert, and a style predictor.
 The denoising network computes face motion conditioned on the speech and style reference video.
@@
 The face motion is rendered into video frames by a renderer [91].
-The style-aware lip expert provides lip motion guidance under diverse expressions and thus drives the denoising network to achieve accurate lip-sync while ensuring style expressiveness.
+The style-aware lip expert provides lip motion guidance under diverse expressions and thus drives the denoising network to achieve accurate lip-sync while preserving emotion fullness.
 The style predictor can predict the speaking style aligned with that conveyed in speech.
@@
 $\bm{m}^{*}_{(0)}=E_{\theta}(\bm{m}_{(t)},t,\bm{A}_{w},\bm{R}).$ (2)
-The asterisk(***) indicates quantities that are generated.
+The asterisk(*) indicates quantities that are generated.
 Our denoising network has a transformer architecture [93].
@@
 The style-aware lip expert encodes the lip motions and audio into respective embeddings conditioned on style reference and then computes the cosine similarity to represent the sync probability.
-To obtain lip motion information from face motion $\bm{m}$, we first convert $\bm{m}$ into the corresponding face mesh and select vertices in the mouth area as the representation of the lip motion [46].
+To obtain lip motion information from face motion $\bm{m}$, we first convert $\bm{m}$ into the corresponding face mesh and select vertices in the mouth area as the representation of the lip motion.
 The lip motion and audio encoders are mainly implemented by MLPs and 1D-convolutions, respectively.
-The style condition is fused into embeddings by first extracting style features from style reference using a style encoder, which mirrors the architecture of the one in the denoising network, and then concatenating the style features with intermediate feature maps from embedding encoders.
-The style encoder in the lip expert and the generator do not share parameters.
+The style condition is fused into embeddings by first extracting style features from style reference using a style encoder, which mirrors the architecture of the one in the denoising network but does not share parameters with it, and then concatenating the style features with intermediate feature maps from embedding encoders.
 Style Predictor.
 Specifically, the style predictor $S_{\phi}$ predicts the style code $\bm{s}$ extracted by the style encoder in the trained denoising network.
-Observing the correlation between speaker identity and style codes (Sec.
-4.4), the style predictor also integrates the portrait as input.
+Since we observe that style codes correlate with speaker identity (section IV-D), the style predictor also integrates the portrait as input.
 The style predictor is instantiated as a diffusion model and is trained to predict the style code itself:
@@
 To obtain the speaker info embedding, our method first extracts the 3DMM identity parameters, which include the face shape information but removes irrelevant information, such as expressions, from the portrait, and then embeds it into a token using an MLP.
+Discussion: Advantages over StyleTalk.
+Although StyleTalk, a GAN-based baseline, also leverages transformer modules, DreamTalk presents notable advantages: 1) StyleTalk’s modules and loss functions are overly complex, which may cause unstable generation results, while DreamTalk’s are simple, making it more extensible and robust.
+Since GAN’s mode-collapse issue hampers modeling diverse speaking styles, to enhance emotion intensity, StyleTalk uses a complex dynamic network and up to six loss terms.
+As discussed in section IV-B, the overly complex design may cause unstable and inferior results.
+In contrast, DreamTalk, leveraging the power of diffusion models, does not need complex modules and only uses two loss terms, which is much simpler.
+2) StyleTalk makes incorrect assumptions about the data, which may impair the performance.
+To apply losses that enhance emotion intensity, StyleTalk assumes the speaking styles are consistent in a predefined video group.
+However, as discussed in section IV-D, these speaking styles are actually varied.
+DreamTalk does not need such an assumption.
+3) StyleTalk can only specify speaking styles using videos, while DreamTalk, leveraging the style predictor, can specify styles only using input audio, which is more convenient.
+TABLE I: Quantitative Comparisons. We do not receive GC-AVT samples on Voxceleb2. SA is only evaluated on MEAD for emotional methods.
 Methods MEAD / HDTF / Voxceleb2
 […]
-MakeItTalk [105] 0.73 / 0.57 / 0.52 0.11 / 0.24 / 0.24 3.97 / 5.12 / 6.29 5.32 / 4.55 / 5.15 2.10 / 3.16 / 2.17
-Wav2Lip [49] 0.80 / 0.63 / 0.54 0.18 / 0.30 / 0.30 2.72 / 4.53 / 5.85 4.05 / 3.60 / 4.64 5.26 / 5.83 / 5.70
-PC-AVS [104] 0.50 / 0.42 / 0.36 0.07 / 0.13 / 0.09 5.83 / 9.71 / 12.9 4.97 / 4.17 / 7.42 2.18 / 4.85 / 4.73
-AVCT [84] 0.83 / 0.76 / 0.64 0.14 / 0.22 / 0.23 2.92 / 2.86 / 3.62 5.52 / 3.57 / 3.71 2.53 / 4.27 / 3.89
-GC-AVT [37] 0.34 / 0.36 /    - 0.14 / 0.28 /    - 8.04 / 10.2 /    - 7.10 / 6.23 /    - 2.42 / 4.72 /    -
-EAMM [30] 0.40 / 0.40 / 0.43 0.08 / 0.14 / 0.20 6.70 / 7.03 / 6.36 6.48 / 6.86 / 4.89 1.41 / 2.54 / 2.24
-StyleTalk [46] 0.84 / 0.81 / 0.66 0.16 / 0.30 / 0.29 2.12 / 1.96 / 2.92 3.25 / 2.41 / 2.96 3.47 / 4.82 / 4.51
-SadTalker [100] 0.69 / 0.77 / 0.44 0.16 / 0.24 / 0.19 4.12 / 5.99 / 9.12 4.37 / 4.07 / 6.11 2.76 / 4.35 / 4.38
-PD-FGC [78] 0.49 / 0.41 / 0.35 0.05 / 0.13 / 0.12 5.50 / 9.50 / 12.5 4.10 / 4.23 / 8.19 2.27 / 4.68 / 4.64
-EAT [20] 0.53 / 0.59 / 0.47 0.15 / 0.26 / 0.20 5.54 / 3.86 / 5.53 4.79 / 4.03 / 5.88 2.16 / 4.54 / 4.35
-DreamTalk 0.86 / 0.85 / 0.69 0.16 / 0.31 / 0.30 1.93 / 1.80 / 2.69 2.91 / 2.15 / 2.72 3.78 / 5.17 / 4.90
-Ground Truth 1 / 1 / 1 0.22 / 0.31 / 0.33 0 / 0 / 0 0 / 0 / 0 4.13 / 5.44 / 5.23
-Table 1: Quantitative comparisons on MEAD, HDTF, and Voxceleb2. Since we only receive GC-AVT samples on MEAD and HDTF, GC-AVT is not evaluated on Voxceleb2.
-3.3 Training and Inference
+MakeItTalk [55] 0.73/0.57/0.52 0.11/0.24/0.24 3.97/5.12/6.29 5.32/4.55/5.15 2.10/3.16/2.17 -
+Wav2Lip [54] 0.80/0.63/0.54 0.18/0.30/0.30 2.72/4.53/5.85 4.05/3.60/4.64 5.26/5.83/5.70 -
+PC-AVS [59] 0.50/0.42/0.36 0.07/0.13/0.09 5.83/9.71/12.9 4.97/4.17/7.42 2.18/4.85/4.73 -
+AVCT [47] 0.83/0.76/0.64 0.14/0.22/0.23 2.92/2.86/3.62 5.52/3.57/3.71 2.53/4.27/3.89 -
+GC-AVT [3] 0.34/0.36/   - 0.14/0.28/   - 8.04/10.2/   - 7.10/6.23/   - 2.42/4.72/   - 18.7
+EAMM [5] 0.40/0.40/0.43 0.08/0.14/0.20 6.70/7.03/6.36 6.48/6.86/4.89 1.41/2.54/2.24 20.1
+StyleTalk [2] 0.84/0.81/0.66 0.16/0.30/0.29 2.12/1.96/2.92 3.25/2.41/2.96 3.47/4.82/4.51 74.2
+SadTalker [60] 0.69/0.77/0.44 0.16/0.24/0.19 4.12/5.99/9.12 4.37/4.07/6.11 2.76/4.35/4.38 -
+PD-FGC [1] 0.49/0.41/0.35 0.05/0.13/0.12 5.50/9.50/12.5 4.10/4.23/8.19 2.27/4.68/4.64 43.9
+EAT [6] 0.53/0.59/0.47 0.15/0.26/0.20 5.54/3.86/5.53 4.79/4.03/5.88 2.16/4.54/4.35 36.4
+DreamTalk 0.86/0.85/0.69 0.16/0.31/0.30 1.93/1.80/2.69 2.91/2.15/2.72 3.78/5.17/4.90 86.7
+Ground Truth 1/1/1 0.22/0.31/0.33 0/0/0 0/0/0 4.13/5.44/5.23 92.5
+III-C Training and Inference
 Training.
 The style-aware lip expert is first pre-trained by determining whether randomly sampled audio and lip motion clips are synchronous as in [54] and then frozen during training the denoising network.
+We use cosine-similarity with binary cross-entropy loss during training.
+Specifically, we compute cosine-similarity for the face motion embedding $\bm{e}^{m}$ and audio embedding $\bm{e}^{a}$ to represent the probability that the input audio-motion pair is synchronized.
+The training loss of the lip expert is:
+$\mathcal{L}_{{\text{expert}}}={\text{BCE}}(\frac{\bm{e}^{m}\cdot\bm{e}^{a}}{\max(||\bm{e}^{m}||_{2}\cdot||\bm{e}^{a}||_{2})}).$ (5)
 The denoising network $E_{\theta}$ is trained by sampling random tuples $(\bm{m}_{(0)},t,\bm{A}_{w},\bm{R})$ from dataset, corrupting $\bm{m}_{(0)}$ into $\bm{m}_{(t)}$ by adding Gaussian noises, executing denoising steps to $\bm{m}_{(t)}$, and optimizing the loss:
@@
 $+(1-\omega)E_{\theta}(\bm{m}_{(t)},t,\bm{A}_{w},\varnothing),$
-instead of Equation 2. This approach enables controlling the effectiveness of the style reference $\bm{R}$ through adjustment of the scale factor $\omega$.
+instead of eq. 2. This approach enables controlling the effectiveness of the style reference $\bm{R}$ through adjustment of the scale factor $\omega$.
 When training the style predictor, we draw a random video, then extract audio $\bm{A}$ and style code $\bm{s}_{(0)}$ (using the trained style encoder) from it.
@@
 $\mathcal{L}_{{\text{pred}}}=\lVert\bm{s}_{(0)}-S_{\phi}(\bm{s}_{(t)},t,\bm{A},\bm{I})\rVert_{2}^{2},$ (10)
-We utilize PIRenderer [52] as the renderer and meticulously fine-tune it to empower the renderer with emotional expression generation capabilities.
+We utilize PIRenderer [91] as the renderer and fine-tune it on emotional dataset to enable it to generate emotional talking faces.
 Inference.
-Our method enables the specification of speaking styles using either reference videos or solely through input audio and portrait.
+Our method can specify speaking styles using either reference videos or solely through input audio and portrait.
 In the case of reference videos, style codes are derived using the style encoder in the denoising network.
-When relying solely on input audio and portrait, these inputs are processed by the style predictor, which employs a denoising procedure to obtain the style code.
+When relying solely on input audio and portrait, the style predictor takes these inputs and employs a denoising procedure to obtain the style code.
 With the style code, the denoising network utilizes the sampling algorithm of DDPM [10] to produce face motions.
@@
 The sampling process can be accelerated by leveraging DDIM [96].
-The output face motions are then rendered into videos by the renderer PIRenderer.
-4 Experiments
-4.1 Experimental Setup
+The output face motions are then rendered into videos by the renderer.
+IV Experiments
+Figure 3: Qualitative comparisons. The red arrow indicates inaccurate lip motions.
+IV-A Experimental Setup
 Datasets.
@@
 Baselines.
-We compare our method with previous methods including: MakeitTalk [105], Wav2Lip [49], PC-AVS [104], AVCT [84], GC-AVT [37], EAMM [30], StyleTalk [46], DiffTalk [58], SadTalker [100], PD-FGC [78], and EAT [20].
-For DiffTalk, since the released model is incomplete and unable to generate reasonable results until submission, we perform qualitative comparisons using videos from its released demo.
-For other methods, we generate the samples using released models or with the help of the authors.
+We compare our method with previous methods: MakeitTalk [55], Wav2Lip [54], PC-AVS [59], AVCT [47], GC-AVT [3], EAMM [5], StyleTalk [2], DiffTalk [18], SadTalker [60], PD-FGC [1], and EAT [6].
+DiffTalk’s released model cannot generate reasonable results until submission, so we perform qualitative comparisons using videos from its demo.
+For other methods, we generate the samples using released models or with authors’ help.
+When generating samples, we use the audio and the first image from the test video as inputs.
+We use a segment of the test video as the reference.
+Except when evaluating the style predictor, the style of DreamTalk is specified by the video.
 Metrics.
-We utilize widely used metrics: SSIM [88], the Cumulative Probability of Blur Detection (CPBD) [48], the SyncNet confidence score ($\text{Sync}_{\text{conf}}$[9], the Landmark Distance around mouth area (M-LMD) [6], the Landmark Distance on the full face (F-LMD).
-Figure 3: Qualitative comparisons with previous methods.
-4.2 Main Results
+To evaluate video quality, we use SSIM [99] and the CPBD [100].
+To evaluate lip-motion accuracy, we use the SyncNet confidence score ($\text{Sync}_{\text{conf}}$[101] and the Landmark Distance around mouth area (M-LMD) [53].
+To evaluate the accuracy of generated expressions, we use the Landmark Distance on the full face (F-LMD) and a newly proposed metric Style Accuracy (SA).
+SA is the accuracy obtained from classifying samples using a speaking style classifier.
+When training the classifier, we divide the MEAD dataset into several groups with approximately consistent speaking styles and train the classifier to sort videos into the correct groups.
+Therefore, if a method generates accurate expressions, its samples will be classified into correct group and hence it will get higher SA.
+The details of SA metric are reported in Supp.
+Mat..
+IV-B Main Results
 Quanitative Comparisons.
-As shown in Tab.
-1, our method outperforms previous methods across most metrics.
-Wav2Lip’s training with SyncNet as a discriminator explains its high SyncNet confidence score, even surpassing the ground truth.
+As shown in table I, our method outperforms previous methods across most metrics.
+Wav2Lip’s SyncNet confidence score is higher than ours, even surpassing the ground truth.
+This is because Wav2Lip is trained using SyncNet as a discriminator.
 Notably, our method’s SyncNet confidence score closely aligns with the ground truth, and it achieves the best M-LMD scores, which indicates its capability for precise lip synchronization.
-Furthermore, our superior performance in the F-LMD metric demonstrates our method’s proficiency in generating facial expressions consistent with the reference speaking style.
-Qualitative comparisons.
-Fig.
-3 shows the qualitative comparisons.
-The portraits, style references, and audio are all unseen during training.
+Furthermore, our superior performance in the F-LMD and SA metrics demonstrates our method’s proficiency in generating facial expressions consistent with the reference speaking style.
+Qualitative comparisons. fig. 3 shows the qualitative comparisons. The portraits, style references, and audio are all unseen during training.
+Figure 4: Comparisons with StyleTalk on in-the-wild style reference. StyleTalk fails to generate accurate emotion.
+Figure 5: Sudden face distortion (marked in red box) that frequently occurs in StyleTalk’s output. Better viewed in Supp. Video
+TABLE II: Comparisons of identity preservation on MEAD. DreamTalk’s score is competitive to non-emotional methods and is the best in emotional methods.
+Method MakeitTalk Wav2Lip PC-AVS AVCT SadTalker GCAVT EAMM StyleTalk PD-FGC EAT TH-PAD Ours Ground Truth
+CSIM $\uparrow$ 0.77 0.89 0.17 0.79 0.78 0.20 0.36 0.64 0.35 0.61 0.30 0.71 0.80
+StyleTalk, one of the most competitive baselines, fails to consistently generate high-quality results across diverse speaking styles.
+Firstly, StyleTalk frequently generates videos with sudden facial distortions (fig.
+5), making it unsuitable for large-scale practical applications.
+We do not observe such phenomena in DreamTalk’s results.
+We speculate that the reason for Styletalk’s unstable generation is that StyleTalk uses overly complex loss functions and is based on GANs, models that suffer from unstable training.
+Secondly, StyleTalk fails to generate rich emotions, especially for in-the-wild style references.
+As shown in fig.
+3, the output from StyleTalk exhibits discrepancies with the style reference: the left speaker’s eyes are not as narrowed, and the right speaker’s mouth is not opened as widely.
+The discrepancies are more pronounced when using in-the-wild style references.
+As shown in fig.
+5, Styletalk fails to generate expressions consistent with those in the style reference, including raised eyebrows, glaring eyes, and widened mouths.
+Initially, we speculate that these issues are caused by the limited training data of StyleTalk, but we observe that when trained only using the same data as StyleTalk, DreamTalk can still generate expressions consistent with the style reference (the result is also shown in fig.
+5).
+Therefore, we speculate that the reason is that GAN’s mode-collapse issue impairs the performance across diverse speaking styles.
+Thirdly, StyleTalk’s lip-sync is inferior.
+A notable example is shown in fig.
+3: when the speaker utters "m"; the expected closed-mouth motion is replaced by an open mouth in StyleTalk’s output.
 It can be seen that MakeItTalk and AVCT struggle with accurate lip synchronization.
@@
 SadTalker, on the other hand, generally aligns lip movements with audio but occasionally displays unnatural jitters.
-EAT’s capability is limited to generating discrete emotions, lacking the finesse for nuanced expressions.
+EAT can only generate discrete emotions, lacking the finesse for nuanced expressions.
 For example, in the left case, the style reference shows the speaker narrowing his eyes, but EAT merely produces a generic disgusted look with wide-open glaring eyes.
 Additionally, as shown in the right case, EAT struggles to maintain a consistent face shape during speaker head movements.
-EAMM, GC-AVT, StyleTalk, and PD-FGC demonstrate the ability to produce fine-grained expressions.
+EAMM, GC-AVT, PD-FGC can produce fine-grained emotions.
 However, EAMM falls short in lip synchronization, GC-AVT and PD-FGC struggle with preserving speaker identity, and all three have issues rendering a plausible background.
-We observed that StyleTalk, while capable of generating nuanced expressions, occasionally does so with diminished intensity and fails to generate accurate lip motion for some words.
-A notable example is shown in the third column of the left case: when the speaker utters "um"; the expected closed-mouth motion is replaced by an open mouth in StyleTalk’s output.
-Besides, as shown in Fig.
-4, StyleTalk occasionally generates videos where the face distorts suddenly.
-Fig. 5 present comparisons with DiffTalk. DiffTalk struggles with lip synchronization and introduces jitteriness and artifacts in the mouth region.
+As shown in fig.
+7, DiffTalk struggles with lip synchronization and produces jitteriness and artifacts in the mouth area.
 In contrast, DreamTalk excels in producing realistic talking faces that not only mirror the reference speaking style but also achieve precise lip synchronization and superior video quality.
-Figure 4: Face distortion observed in StyleTalk’s output.
-Figure 5: Comparisons with DiffTalk.
-Figure 6: The results generated using out-of-domain portraits.
+Figure 6: Comparisons with DiffTalk. DiffTalk fails to achieve lip-sync and produces jitteriness.
+Figure 7: Emotional results generated from songs in multiple languages (English, Chinese, Japanese).
+Figure 8: The results of speaking style prediction. The fourth column displays samples generated with predicted styles applied to the same portrait for clearer comparisons.
+Evaluation of the Identity Preservation.
+To evaluate the ability to preserve identity, we utilize a widely used metric CSIM.
+When computing the CSIM score, we utilize an off-the-shelf face recognition network ArcFace [102] to extract the deep identity features from each generated frame and then calculate the cosine similarities between the features of the input portrait and the generated frames.
+We find that when the ID in the style reference and portrait differ significantly, ID preservation worsens.
+So, we compute scores for StyleTalk and DreamTalk when the style reference comes from a randomly selected ID different from that in portrait.
+table II shows the result on MEAD dataset.
+Wav2Lip attains the highest score since it merely changes the mouth region and leaves other parts intact.
+Non-emotional methods achieve better scores than emotional ones because changing emotions may change the identity perceived by humans and ArcFace.
+DreamTalk’s score is competitive to non-emotional methods and is the best in emotional methods.
 Generalization Capabilities.
-Supplementary Video demonstrates our method’s capability to produce realistic talking head videos for out-of-domain portraits (paintings, toonified portraits, and sculptures, shown in Fig.
-6), speech in various languages, noisy audio input, and songs.
-Sec.
-B.1 provides more analysis.
-Figure 7: The results of speaking style prediction. The fourth column displays samples generated with predicted styles applied to the same portrait for clearer comparisons.
+Leveraging the strong power of diffusion models, DreamTalk shows strong generalization capabilities.
+As shown in fig.
+7 and Supp.
+Video, DreamTalk can even generate reasonable results for songs in multiple languages, even though our training dataset includes only a few multilingual audio and no song audio.
+Supp.
+Video shows that DreamTalk also generalizes well to speech in various languages and noisy audio.
 Results of Speaking Style Prediction.
@@
 8 presents the results of speaking style predictions.
-The speakers with angry emotion are sampled from MEAD while those with fear emotion are sampled from RAVEDESS.
-The style predictor, utilizing emotional audio and neutral portraits from test videos, adeptly deduces personalized speaking styles that align with those observed in the original videos.
-It demonstrates the capacity to discern subtle expressions within the same emotion.
+The style predictor, utilizing emotional audio and neutral portraits, adeptly deduces personalized speaking styles aligned with those in the original videos.
+It can discern subtle expressions within the same emotion.
 For instance, for samples with angry emotion, the first-row speaker exhibits narrowed eyes, in contrast to the second-row speaker’s intense, glaring stare.
-For samples with angry emotion, the first-row speaker’s eyes and mouth are open, whereas the second-row speaker combines narrowed eyes with a contorted facial expression.
-The ablation study and user study for the style predictor are presented in Appendix A.
-4.3 Ablation Study
-Figure 8: Ablation study results
+For samples with fear emotion, the first-row speaker’s eyes and mouth are open, whereas the second-row speaker combines narrowed eyes with a contorted facial expression.
+Figure 9: Analyzing the influence of portraits on style prediction. The audio conveys surprised emotion.
+We analyze the influence of portraits on speaking style prediction by predicting speaking styles with an audio clip and different input portraits.
+The predicted styles are subsequently applied to an identical portrait for a clearer comparison.
+As shown in fig.
+9, the predicted speaking styles match the subtle identity characteristics, such as gender, of the input portraits.
+The predicted style A generated more feminine results.
+This validates the necessity of integrating portrait information during style prediction.
+Figure 10: Comparisons with TH-PAD.
+Figure 11: Ablation study results of emotional talking head generation.
+We compare the style predictor with TH-PAD [19], a method that also uses audio to predict expressions.
+We obtain TH-PAD samples with the authors’ help.
+For comparisons, we use audio to predict speaking styles and use predicted styles to generate samples.
+We observe that TH-PAD fails to generate emotions conveyed in audio.
+TH-PAD only generates neutral expressions aligned with the audio rhythm.
+fig.
+11 shows that TH-PAD fails to generate the sad emotion reflected in the audio.
+We also conduct a quantitative comparison on MEAD.
+As shown in table III, TH-PAD’s SA and LMD scores, which measure emotion alignment, are inferior.
+TABLE III: Comparisons with TH-PAD on MEAD. DreamTalk(A) uses styles predicted from audio.
+[…]
+       TH-PAD        6.38/5.81        5.3
+       DreamTalk(A)        2.24/3.43        78.6
+       Ground Truth        0 / 0        92.5
+TABLE IV: The results of DreamTalk’s ablation study on MEAD. CPBD is omitted due to no significant differences.
 […]
@@
 Full 0.86 1.93 2.91 3.78
-Table 2: The results of DreamTalk’s ablation study on MEAD. We omit CPBD scores since there are no significant differences between variants on CPBD.
+IV-C Ablation Study
+Emotional Talking Head Generation.
 To analyze the contributions of our designs, we conduct an ablation study with two variants: (1) remove the style-aware lip expert (w/o lip expert); (2) trained with unconditional lip expert (uncond lip expert).
@@
 fig.
-8 and Tab.
-2 present our ablation study results.
+11 and table IV present our ablation study results.
 The variant w/o lip expert exhibits a decline in lip-sync accuracy on the emotional dataset MEAD, despite its competitive F-LMD score indicating expressive facial generation.
@@
 The Full model achieves a harmonious balance, ensuring both precise lip synchronization and vivid expressions, thanks to the style-aware lip expert directing the diffusion model’s expressive potential.
-4.4 Style Code Visualization
-Figure 9: t-SNE visualization of style codes from 15 speakers. Each color stands for style codes from an identical speaker.
-Using t-distributed stochastic neighbor embedding (t-SNE) [75], we map style codes from the MEAD dataset’s 15 speakers into a 2D space.
-These speakers exhibit 22 distinct speaking styles, comprising seven emotions at three intensity levels, alongside a neutral style.
-For each style, style codes are extracted from 10 randomly chosen videos.
-Fig.
-9 reveals that style codes from identical speakers tend to cluster, despite some codes sharing identical emotion categories.
-This suggests that the variance in speaking styles attributable to individual speaker idiosyncrasies outweighs that induced by different emotions, thereby underpinning the rationale for using portrait information to infer speaking styles.
-We also observe that each speaker’s style code distribution exhibits both common patterns and individualized characteristics.
-We present these intriguing observations in Sec.
-B.2.
-4.5 Style Manipulation
-Figure 10: Effect of the input style controlled by adjusting the scale $\omega$ of the classifier-free guidance.
-Adjusting the scale of Classifer-free Guidance.
-As elaborated in Sec.
-3.3, the scale factor $\omega$ in the classifier-free guidance scheme modulates the effect of the input style.
-Adjusting $\omega$ either amplifies or attenuates the designated style, as shown in Fig.
-10.
-When $\omega=0$, DreamTalk produces a talking head with a neutral expression.
-We observed that when the scale factor $\omega$ exceeds $2$, there is a noticeable decline in lip-sync accuracy.
-Figure 11: The results of speaking style interpolation.
-Style Code Interpolation.
-Leveraging the style space, we can modify speaking styles via style code manipulation.
-Fig.
-11 illustrates that linear interpolation between style codes results in a seamless transition of generated speaking styles.
-This interpolation process allows for style intensity modulation and the generation of novel speaking styles.
-4.6 User Study
-[…]
-MakeItTalk [105] 1.94 2.03 1.65
-Wav2Lip [49] 2.29 1.45 1.14
-PC-AVS [104] 2.26 1.81 1.86
-AVCT [84] 2.31 2.21 1.72
-GC-AVT [37] 2.35 1.20 1.88
-EAMM [30] 1.81 1.40 1.78
-StyleTalk [46] 2.35 2.29 2.08
-SadTalker [100] 2.37 2.38 1.73
-PD-FGC [78] 1.95 1.61 2.31
-EAT [20] 2.24 1.65 2.29
-DreamTalk 2.55 2.60 2.46
-Ground Truth 3.03 2.89 1.83
-Table 3: User study results.
-We conduct a user study of 20 participants.
-We generate the test samples covering multiple speaking styles and speakers.
-For each method, the participant is required to score $10$ videos sampled from the test samples and is asked to give a rating (from 1 to 5, 5 is the best) on three aspects: (1) the Lip sync quality, (2) the realness of results, and (3) the style consistency between the generated videos and the style reference.
-As shown in Tab.
-3, our method outperforms existing approaches across all aspects, particularly in style consistency, highlighting its superior capabilities.
-5 Conclusion
-In this work, we propose DreamTalk, a novel approach leveraging diffusion models for generating expressive talking heads.
-Our method aims to excel in diverse speaking styles while minimizing dependence on extra style references.
-We develop a denoising network for creating expressive, audio-driven facial motions and introduce a style-aware lip expert to optimize lip-sync without compromising style expressiveness.
-Additionally, we devise a style predictor that infers speaking styles directly from audio, eliminating the need for video references.
-The efficacy of DreamTalk is validated through extensive experiments.
-Acknowledgements.
-This work is supported by Alibaba Group through Alibaba Research Intern Program.
-We would like to thank Xinya Ji, Borong Liang, Yan Pan, and Suzhen Wang for their generous help with the comparisons.
-Supplementary Material
-Appendix A Additional Evaluation for Speaking Style Prediction
-A.1 Ablation Study
 Figure 12: The qualitative results of style predictor’s ablation study.
+Speaking Style Prediction.
 To evaluate the impact of our design choices, we conduct an ablation study with three variants: (1) omitting speaker information and relying solely on audio for prediction (w/o speaker info); (2) during model training, the speaker info and audio are both obtained from the same video (w/o cross-ID training); (3) employing a regression model instead of a diffusion model for prediction (regression).
 Our full model is denoted as Full.
-When generating samples for evaluation, the facial images and audio we use are sourced from videos of the same individual expressing different emotions(e.g.
+When generating samples for evaluation, the facial images and audio we use are sourced from videos of the same individual expressing different emotions(e.
+g.
 the face image is from a happy video while the audio is from an angry one.).
@@
   • Motion Distance (MD) We use the predicted style codes and the audio used for prediction to generate face motions and compute the L2 distance between the generated face motions and the face motions extracted from the ground truth videos.
-  • Style Accuracy (SA) We split test videos into different speaking styles and train a style classifier to classify which style a face motion sequence belongs to.
-Then, We classify the face motions generated using predicted style codes and report the accuracy.
-Specifically, we put the videos from the same speaker, emotion, and intensity into one style.
-We evaluate this metric on MEAD only since the number of RAVEDESS videos for each style is inadequate to train a style classifier.
-The ground truth testing set gets $92.5\%$ accuracy.
+  • Style Accuracy (SA) We use the SA metric mentioned in the section IV-A. SA is evaluated on 3DMM face motions. The ground truth testing set gets $92.5\%$ accuracy.
 We refrain from devising image-level metrics, such as training an image classifier for speaking style classification, due to several critical considerations.
@@
 Secondly, inaccuracies introduced by the rendering process may further additionally hinder the accurate discernment of these subtle speaking styles.
+TABLE V: The ablation study results of the style predictor.
 […]
@@
 Full 0.42 0.23 78.6
-Table 4: The ablation study results of the style predictor.
-The results are shown in Tab.
-4 and Fig.
+The results are shown in table V and fig.
 12.
@@
 The regression variant struggles to generate accurate expressions for certain data, highlighting the superior distribution-learning capability of diffusion models in facilitating speaking style prediction.
-A.2 User Study
+IV-D Style Code Visualization
+Figure 13: t-SNE visualization of style codes.
+Left: Style codes from 15 speakers.
+Each color indicates style codes from an identical speaker.
+Right: Style codes from a speaker, with darker hues representing increased emotional intensity.
+Figure 14: Modulating style intensity by adjusting the scale $\omega$ of classifier-free guidance.
+We use t-SNE [103] to map style codes from the MEAD dataset’s 15 speakers into a 2D space.
+Each speaker exhibits 22 distinct speaking styles, comprising seven emotions at three intensity levels, alongside a neutral style.
+For each style of each speaker, we randomly select 10 videos to extract style codes for visualization.
+The left in fig.
+13 shows that style codes first cluster based on identity rather than emotion, indicating the correlation between speaker identity and speaking styles, hence justifying the rationale for using portrait information to infer speaking styles.
+The right in fig.
+13 shows that even within the same emotion and the same intensity, a speaker’s expressions can vary significantly.
+The speaker expresses intense sadness in two ways.
+In one way, the speaker clenches teeth (top left portrait), similar to happy expressions (bottom left portrait), while in another way, the speaker depresses lip corners (bottom right portrait).
+The observation suggests StyleTalk’s assumption that speaking styles are consistent in videos with the same emotions is incorrect, which may impair performance.
+IV-E Modulating Style Intensity
+As elaborated in section III-C, the scale factor $\omega$ in the classifier-free guidance scheme can modulate the intensity of any input style, even those unseen during training.
+As shown in fig.
+14, the style reference is in-the-wild and adjusting $\omega$ either amplifies or attenuates the designated style.
+When $\omega=0$, DreamTalk produces a talking head with a neutral expression.
+We observed a noticeable decline in lip-sync accuracy when the scale factor $\omega$ exceeds $2$.
+IV-F User Study
+TABLE VI: Mean ratings of user study with $95\%$ confidence intervals.
+[…]
+MakeItTalk [55] 2.73 $\pm$ 0.08 3.10 $\pm$ 0.08 -
+Wav2Lip [54] 3.46 $\pm$ 0.07 2.45 $\pm$ 0.05 -
+PC-AVS [59] 3.29 $\pm$ 0.04 2.86 $\pm$ 0.07 -
+AVCT [47] 3.43 $\pm$ 0.07 3.39 $\pm$ 0.08 -
+GC-AVT [3] 3.40 $\pm$ 0.06 2.21 $\pm$ 0.05 2.70 $\pm$ 0.05
+EAMM [5] 2.83 $\pm$ 0.08 2.46 $\pm$ 0.07 2.43 $\pm$ 0.06
+StyleTalk [2] 3.42 $\pm$ 0.05 3.48 $\pm$ 0.07 2.88 $\pm$ 0.07
+SadTalker [60] 3.44 $\pm$ 0.07 3.51 $\pm$ 0.06 -
+PD-FGC [1] 3.16 $\pm$ 0.06 2.80 $\pm$ 0.06 3.06 $\pm$ 0.07
+EAT [6] 3.39 $\pm$ 0.07 2.87 $\pm$ 0.08 2.91 $\pm$ 0.07
+DreamTalk 3.80 $\pm$ 0.07 3.77 $\pm$ 0.07 3.34 $\pm$ 0.08
+Ground Truth 4.21 $\pm$ 0.08 4.09 $\pm$ 0.08 -
+Emotional Talking Head Generation.
+We conduct a user study to further evaluate our method.
+We generate 30 test samples for each method, which cover diverse speaking styles and speakers, and recruit 22 participants to rate samples.
+Each participant is required to rate all samples (from 1 to 5, 5 is the best) on three aspects: (1) lip sync quality, (2) video realness, and (3) style consistency between the generated videos and the style reference (This metric is only evaluated on emotional methods.
+Since Ground Truth videos do not express the expressions reflected in the style references.
+The score for Ground Truth is omitted.).
+As shown in table VI, our method outperforms baselines across all aspects.
+A one-way ANOVA and a post-hoc Tukey test identify a significant difference ($p<0.001$) between our method and other baselines on all aspects.
+Speaking Style Prediction.
 In our user study, we evaluate the alignment between the original and predicted speaking styles.
@@
 This indicates that the style predictor is able to infer personalized speaking styles that are aligned with the audio.
-A.3 Analyzing the Influence of Portraits
-Figure 13: Analyzing the influence of portraits on style prediction. The audio conveys surprised emotion.
-We analyze the influence of portraits on speaking style prediction by predicting speaking styles with an audio clip and different input portraits.
-The predicted styles are subsequently applied to an identical portrait for a clearer comparison.
-As shown in Fig.
-13, the predicted speaking styles match the subtle identity characteristics, such as gender, of the input portraits.
-The predicted style A generated more feminine results.
-This validates the necessity of integrating portrait information during style prediction.
-Appendix B Additional Results for Expressive Talking Head Generation
-B.1 Analysis on Generalization Capabilities
-Songs.
-As demonstrated in Supplementary Video, our method successfully generates reasonable results for songs, even those with accompaniment, despite this being significantly different from the training dataset’s data distribution.
-A noticeable decline in lip-sync accuracy is observed when the accompaniment volume is excessively high.
-We conduct a comparative analysis of lip-sync performance between songs with accompaniment and songs with removed accompaniment (using data from the SingFace Dataset [38]).
-It is found that the accompaniment adversely affects lip-sync, leading to mouth movements resembling mumbling.
-Addressing the negative impact of accompaniment on lip-sync accuracy presents an interesting avenue for future research.
-Speech in Multiple Languages.
-Supplementary Video shows that our method generates satisfactory results with speech in French, Chinese, Spanish, German, Italian, Japanese, and Korean.
-The versatility of wav2vec features aids application across various languages.
-Additionally, the inclusion of multilingual talking head videos from Voxceleb2 enhances generalization.
-Noisy Audio.
-Supplementary Video demonstrates that our method yields satisfactory outcomes when processing audio mixed with multiple noise types and intensities.
-We employ noise recordings from typical talking head application environments—meetings, offices, and cafeterias—sourced from the DEMAND dataset [73].
-Using a publicly available tool [^1] https://github.com/Sato-Kunihiko/audio-SNR, we blend the audio with noise at SNRs of 20 dB, 10 dB, and 0 dB.
-Remarkably, our method maintains performance even at 0 dB SNR, where the noise is as loud as the speech and significantly impairs speech intelligibility and clarity.
-B.2 More Results of Style Code Visualization
-Figure 14: t-SNE visualization of style codes for 3 speakers, with darker hues representing increased emotional intensity.
-We observe that each speaker’s style code distribution exhibits both common patterns and individualized characteristics.
-Common patterns include: Firstly, speaking styles of different emotions cluster together first, with styles of lower intensity being closer to neutral and those of higher intensity being further away.
-Secondly, speaking styles of anger and disgust, as well as fear and surprise, often cluster together, as shown in Fig.
-14 (b) and (c).
-Note that unlike Ma et al.
-[46], our method does not incorporate losses to constrain style space.
-Fig.
-14 (a) illustrates an example of individualized characteristics.
-The speaker’s manifestation of fear closely resembles sadness, lacking the characteristic wide-eyed and open-mouthed expression, thereby positioning the speaking styles of fear nearer to those of sadness rather than surprise.
-Even within the same emotion, a speaker’s speaking style can exhibit notable variation.
-In Fig.
-14 (c), the speaker’s dual expression of sadness—once with clenched teeth, similar to happy expressions, and another with depressed lip corners, akin to disgust—results in style codes close to the respective emotions.
-This observation diminishes the rationale for manually categorizing styles based on emotion and intensities in Ma et al.
-[46].
-Appendix C Implementation Details
-C.1 Architectual Details
-Denoising Network.The audio encoder processes an input window of 11 sequential audio features, each of dimension $1024$.
-These features undergo dimension reduction to $256$ via a linear layer and then are fed to a transformer encoder comprising three 8-head transformer encoder layers, each with a hidden size of 256.
-Subsequently, a linear layer transforms the output tokens to yield audio tokens sized $11\times 256$.
-The audio tokens are concatenated with noisy motion and then added with the encoded diffusion step.
-The style encoder ingests sequential expression parameters from style reference videos, each sequence sized $N\times 64$.
-These sequences, ranging in length from 64 to 256 frames, are initially expanded to 256 dimensions via a linear layer.
-Subsequently, they are introduced into a transformer encoder, composed of three 8-head layers, each with a hidden size of 256.
-The resulting output tokens, each with a dimension of $256$, are aggregated through self-attention pooling [56], yielding a style code of dimension $256$.
-Within the decoder, the style code is repeated $11$ times, subsequently added with positional embedding to produce style tokens.
-These tokens, in conjunction with audio tokens, are processed by a transformer decoder, encompassing three 8-head layers, each with a hidden dimension of 256.
-Here, style tokens serve as the query, while audio tokens serve as both key and value.
-The middle output token is fed into a linear layer to predict facial motion.
-Style-aware Lip Expert.
-Figure 15: The architecture of mouth embedder and audio embedder.
-ConvBlock consists of a 1D CNN layer and a batch normalization layer.
-"k3, 32, s2" means that the kernel size is 3, the output dimension is 32, and the stride is 2.
-ConvRBlock is a ConvBlock with residual connection.
-The face mesh is obtained by adding the mean shape to the product of the expression parameters and expression bases. Fig. 15 shows the architecture of the audio embedder and the mouth embedder.
-Style Predictor.
-The style predictor is implemented as a transformer encoder comprising six 8-head transformer encoder layers, each with a hidden size of $256$.
-The input features are all linearly projected to $256$.
-C.2 Data Details
-C.2.1 Datasets
-MEAD.
-The dataset, an in-lab talking-face corpus, features 60 speakers articulating eight emotions at three different intensity levels.
-When dividing the MEAD dataset into training and test subsets, we adhere to previously established methodologies [30].
-HDTF. The dataset stands out as a high-resolution, in-the-wild audio-visual dataset. We designate $10\%$ of HDTF videos for testing and reserve the remainder for training.
-Voxceleb2.
-Voxceleb2 is a large-scale talking head dataset with videos collected from YouTube.
-We redownload and recrop the videos to improve their resolution to $256\times 256$.
-Subsequently, approximately $80000$ high-quality videos are selected, with $400$ allocated for testing and the rest for training.
-RAVDESS.
-The dataset features 24 professional actors (12 female, 12 male) vocalizing two lexically matched statements in a neutral North American accent.
-It encompasses a range of expressions in both speech and song, each articulated at two levels of emotional intensity, with an additional neutral expression included.
-We employ the speech data from RAVDESS, encompassing eight emotions, for evaluation.
-C.2.2 Data Processing
-The original videos are cropped and resized to $256\times 256$ pixels, aligning with the specifications in FOMM [59], and are sampled at 25 FPS. The 3DMM parameters are extracted by Deep3DFace [15].
-Regarding the audio features used in the denoising network and the style-aware lip expert, we downsample the speech wave into the sampling rate of 16000 and extract acoustic features employing a pre-trained Wav2Vec2.0 model [2].
-For audio features used in the style predictor, we extract them using a pre-trained HuBERT model [28].
-Besides, we also utilize low-level audio features including Mel Frequency Cepstrum Coefficients (MFCC), Mel-filterbank energy features (FBANK), fundamental frequency, and voice flag.
-These two type of features are concatenated to represent the audio features used in the style predictor.
-C.3 Training Details
-Our framework is implemented on Pytorch.
-We employ Adam [32] for optimization, with a learning rate set to $0.0001$.
-The number of diffusion steps for the denoising network and style predictor is $1000$.
-The training batch size for the denoising network, style predictor, and lip expert is $64$, $64$, and $32$, respectively.
-$\lambda_{\text{denoise}}$, $\lambda_{\text{sync}}$, $n$, and $w$ are set to $1$, $1$, $5$, and $5$, respectively.
-The number of frames for style reference and audio used in style prediction is limited to $64-256$, corresponding to a time length of $2.56-10.24$ seconds.
-The denoising network, style predictor, and lip expert are trained on one NVIDIA Tesla A100 GPU for about $3$, $1$, and $10.5$ hours, respectively.
-C.3.1 Finetuning PIRender
-The renderer is fine-tuned with the losses in  Ren et al.
-[52] using MEAD.
-Instead of training with the self-reconstruction protocol where the source frame and target frame are from the same video, we select the source frame and target frame from the same speaker with different emotions.
-This enables the renderer to generate emotions different from the input portrait.
-This also allows our method to utilize portraits with emotions, unlike previous approaches that are confined to using neutral portraits [30].
-We also observed that when fine-tuned only on the emotional talking head dataset, the renderer struggles with identity preservation.
-We argue that this problem stems from the fact that the number of speakers in the emotional dataset is limited.
-Therefore, we incorporate some neutral videos in Voxceleb into the data used for fine-tuning.
-This enhances the performance in identity preservation.
-C.3.2 Training Style-aware Lip Expert
-The style-aware lip expert is trained to discriminate whether the input audio and face motions are synchronized.
-We use cosine-similarity with binary cross-entropy loss to train the lip expert.
-Specifically, we compute cosine-similarity for the face motion embedding $\bm{e}^{m}$ and audio embedding $\bm{e}^{a}$ to represent the probability that the input audio-motion pair is synchronized.
-The training loss of the lip expert is:
-$\mathcal{L}_{{\text{expert}}}={\text{BCE}}(\frac{\bm{e}^{m}\cdot\bm{e}^{a}}{\max(||\bm{e}^{m}||_{2}\cdot||\bm{e}^{a}||_{2},\epsilon)}),$ (10)
-where $\epsilon$ is a small number for avoiding the division-by-zero error.
-C.4 Inference Details
-The inference of the denoising network can be accelerated with DDIM.
-We generate samples with $10$ DDIM steps and observe no performance drop.
-Generating a 30-second video offline takes $15.61$ seconds, with the face motion generation only taking $1.24$ seconds.
-During evaluation, the scale factor $\omega$ of classifier-free guidance is set to $1$.
-The style predictor uses the sampling algorithm of DDPM to predict style codes.
-The emotion conveyed in the style reference (video or audio), should remain consistent to avoid confusing the model.
-The head pose information, which is fed into the renderer, can be derived from real videos or generated using existing methods [83].
-Appendix D Limitations and Future Work
-Despite DreamTalk’s promising advancements in expressive talking head generation, it encounters several challenges that open avenues for future research.
-Firstly, the method occasionally produces artifacts, such as teeth flickering, around the mouth area, particularly during intense expressions.
-Generating teeth is a long-standing challenge in talking head generation since the algorithm needs to inpaint the teeth area that are often occluded in the input portrait.
-This problem is exacerbated under intense expressions where the teeth area expands.
-The issue can be mitigated by incorporating modules [44, 92, 36] proposed recently that enhance the teeth quality.
-A more comprehensive solution involves developing an emotion-specific renderer, as current renderers [90, 46, 20, 30] are mainly adaptations of existing facial reenactment methods with minimal modifications.
-An emotion-aware renderer would not only address the teeth generation issue but also enhance the overall expressiveness of emotions.
-Secondly, DreamTalk does not account for the temporal variability in speaking styles.
-In real-life scenarios, a speaker’s style evolves over time, a feature our method currently overlooks.
-For instance, at the end of a speech, our method might still produce expressions of intense emotion, such as a wide-open mouth in surprise, instead of a more neutral, closed-mouth expression.
-Introducing a module that dynamically predicts speaking style over time could address this limitation.
-Thirdly, the style predictor sometimes struggles with accurately identifying emotions in low-intensity audio clips from the MEAD dataset (some audio in MEAD intensity level 1 videos).
-The emotion conveyed in these audio clips is very similar to neutral emotion and thus confusing the prediction.
-We also observed that although the videos in the MEAD dataset clearly express emotions due to the good training and supervision of the speakers, sometimes the audio does not correspond with the emotions that should be expressed.
-Therefore, To enhance prediction accuracy, employing a dataset where the audio closely aligns with the expressed emotions could be beneficial.
+V Limitations
+Despite DreamTalk’s promising advancements in emotional talking head generation, it encounters several challenges that open avenues for future research.
+When using a constant style reference, DreamTalk generates expressions that are strictly consistent with the main expressions in the reference but lack expression changes over time, such as eye blinking.
+Generating emotions with rich temporal variations is, compared to methods that generate neutral or coarse-grained emotions, more difficult for methods that achieve fine-grained emotional control, like DreamTalk.
+This is because these methods must achieve both precise control of expressions and diverse changes.
+To address the issue, temporal changes in expressions can be achieved by using temporally changing style references when generating each video frame.
+As discussed in Supp.
+Mat., DreamTalk can achieve smooth expression changes by changing the style references.
+Eye blinks can be achieved by using eye blink loss [60] during training or post-editing 3DMM, a common practice used in previous methods [104, 66] that also aim to control expressions.
+Specifically, we can obtain the parameter changes of blinking and then edit the generated expression parameters.
+A video with the eye blinking through post-editing is shown in Supp.
+Video.
+DreamTalk may change the speaker’s identity when the identity in the reference video highly differs from that in the portrait.
+The reason is that 3DMM expression parameters leak identity information.
+This leakage leads to the generated identity becoming somewhat similar to the identity in the reference.
+The issue can be mitigated by adopting 3DMM which decouples expression and identity information more effectively.
+The style predictor sometimes struggles with accurately identifying emotions in low-emotion-intensity audio clips from the MEAD dataset.
+The reason is that in some MEAD videos, the audio does not correspond with the expressed emotions.
+To enhance prediction accuracy, it is beneficial to employ a dataset where the audio closely aligns with the expressed emotions.
 Another solution is to incorporate textual information from audio during prediction, a strategy commonly employed in speech emotion recognition [105, 106].
-Despite these challenges, DreamTalk marks a significant stride in the realm of high-quality, expressive talking head generation, setting a foundation for further innovations.
-Appendix E Ethical Consideration
-DreamTalk is able to generate realistic talking head videos.
-This positions DreamTalk with a broad spectrum of potential applications, each carrying intricate societal implications.
-While DreamTalk holds significant potential in amplifying and enriching human creative endeavors and may pave the way for innovative tools for creative professionals, its capabilities also harbor risks.
-There’s a possibility for DreamTalk to generate content that might encompass or imply sexual themes, promote hatred, or depict violence.
-Misuse of DreamTalk could lead to negative repercussions on individuals or groups, potentially erasing or maligning them, perpetuating stereotypes, and subjecting them to disrespect.
-Other concerns include the potential for harassment, intimidation, or exploitation.
-Furthermore, DreamTalk’s capabilities might be harnessed to mislead or spread misinformation.
-Before releasing DreamTalk, we have implemented and plan to introduce several safeguards to curb potential misuse.
-We’ve purged detrimental content from the training dataset and incorporated visual filters to deter users from creating harmful outputs.
-To counteract biases in the generated results, we’re enhancing the dataset’s diversity by manually ensuring balance, which will reduce instances of erasure, stereotype perpetuation, indignity, and uneven quality across inputs.
-Users will be advised against using images without the depicted individuals’ consent to combat harassment and bullying.
-To prevent the spread of misinformation, all DreamTalk outputs will bear watermarks indicating their synthetic nature.
-Our pre-release strategy involves a thorough risk assessment, leveraging a growing suite of safety evaluations and red teaming techniques.
-We’ll also scrutinize the findings from pilot tests centered on new use cases and conduct in-depth post-release assessments.
-Both automated and manual monitoring mechanisms are in development to preempt misuse.
-Our commitment remains steadfast in continuously researching ways to minimize adverse societal effects.
+DreamTalk occasionally produces artifacts around the mouth area, such as teeth flickering, particularly during intense expressions.
+The issue comes from the renderer and can be mitigated by using more advanced renderers.
+Despite these challenges, DreamTalk marks a significant stride in the realm of emotional talking head generation, paving the way for further innovations.
+VI Conclusion
+In this work, we propose DreamTalk, a novel diffusion-based framework that can consistently generate high-quality talking heads in diverse speaking styles and conveniently use audio to specify personalized emotions.
+We develop a denoising network for creating emotional, audio-driven facial motions and introduce a style-aware lip expert to optimize lip-sync while maintaining emotion intensity.
+Additionally, we devise a style predictor that infers speaking styles directly from audio, eliminating the need for video references.
+Extensive experiments validate the efficacy of DreamTalk.
+The results demonstrate that employing diffusion models markedly improves the quality of emotional talking head generation.
+VII Ethical Consideration
+DreamTalk can generate vivid talking head videos, opening up diverse applications but also posing risks like promoting hatred or depicting violence.
+Misuse could harm individuals or groups, perpetuate stereotypes, and spread misinformation.
+To mitigate these risks, we’ve implemented safeguards like watermarks on all outputs and advising against using images without consent.
+We remain committed to continuous research to minimize adverse societal impacts.

### v3

@@
 Although these methods achieve satisfactory results for a limited range of speaking styles, they struggle with more diverse styles, especially ones unseen during training, often resulting in diminished emotional intensity, inaccurate lip motion, or sudden facial distortions [3].
-Another issue is that to specify speaking styles, previous methods often rely on extra references, such as videos [3, 5, 1] or texts [8, 9].
+Another issue is that to specify speaking styles, previous methods often rely on extra references, such as videos [4, 6, 2] or texts [9, 10, 11].
 Their acquisition requires extra manual effort and hence is inconvenient.
@@
 Audio-Driven Talking Head Generation.
-Audio-driven methods [22, 23, 24, 25, 26, 27] fall into two main categories: person-specific and person-agnostic.
+Audio-driven methods [24, 25, 26, 27, 28] fall into two main categories: person-specific and person-agnostic.
 Person-specific approaches [29, 30, 31, 32, 33] are constrained to generating videos for speakers seen during training.
-Many of these [33, 34, 35, 36, 37, 30, 38, 39, 40] first craft 3D facial animations, later converting them into realistic videos.
+Many of these [34, 35, 36, 37, 31, 38, 39, 40] first craft 3D facial animations, later converting them into realistic videos.
 Recent advancements [41, 42, 43, 44, 43, 45] have employed neural radiance fields for modeling, yielding high-fidelity, realistic videos.
@@
 Early methods [10, 7, 62, 63, 64, 32, 31, 65, 66] model expressions in discrete emotions.
-To model fine-grained emotions, recent methods [5, 3, 2, 67] use an expression reference video and transfer the expressions from that video to the generated one.
+To model fine-grained emotions, recent methods [6, 4, 3, 1, 67] use an expression reference video and transfer the expressions from that video to the generated one.
 However, these GAN-based methods cannot consistently achieve high performance across diverse emotions.
@@
 Diffusion models [13, 12] have demonstrated strong performance across multiple vision tasks [14, 16, 71, 72, 73, 74], including text-to-image generation [75], human motion generation [19], and video generation [17, 76, 77, 78, 79].
-Most diffusion-based methods for talking head generation [18, 80, 81, 82, 83, 84, 19, 85, 86, 87, 88], including EMO [21], AniPortrait [85] and Hallo [89], mainly generate talking heads with neutral emotions and lacks emotional controllability.
+Most diffusion-based methods for talking head generation [20, 80, 81, 82, 83, 84, 21, 85, 86, 87, 88, 89], including EMO [23], AniPortrait [85] and Hallo [90], mainly generate talking heads with neutral emotions and lacks emotional controllability.
 Besides, the inference of EMO is slow.
@@
 This leakage leads to the generated identity becoming somewhat similar to the identity in the reference.
-The issue can be mitigated by adopting 3DMM which decouples expression and identity information more effectively.
+The issue can be mitigated by adopting 3DMM which decouples expression and identity more effectively.
 The style predictor sometimes struggles with accurately identifying emotions in low-emotion-intensity audio clips from the MEAD dataset.
@@
 To enhance prediction accuracy, it is beneficial to employ a dataset where the audio closely aligns with the expressed emotions.
-Another solution is to incorporate textual information from audio during prediction, a strategy commonly employed in speech emotion recognition [105, 106].
+Another solution is to incorporate textual information from audio during prediction, a strategy commonly employed in speech emotion recognition [106].
 DreamTalk occasionally produces artifacts around the mouth area, such as teeth flickering, particularly during intense expressions.